inputenc.sty

Alan Jeffrey  Frank Mittelbach

v0.95 1997/12/20

1 Introduction

This package allows the user to specify an input encoding (for example, ASCII, ISO Latin-1 or Macintosh) by saying:

  \usepackage[encoding name]{inputenc}

The encoding can also be selected in the document with:

  \inputencoding{encoding name}

The only practical use of this command within a document is when using text from several documents to build up a composite work such as a volume of journal articles. Therefore this command will be used only in vertical mode.

The encodings provided by this package are:

• ascii ASCII encoding for the range 32–127.

• latin1 ISO Latin-1 encoding.

• latin2 ISO Latin-2 encoding.

• latin3 ISO Latin-3 encoding.

• latin5 ISO Latin-5 encoding.

• decmulti DEC Multinational Character Set encoding.

• cp850 IBM 850 code page.

• cp852 IBM 852 code page.

• cp437 IBM 437 code page.

• cp437de IBM 437 code page (German version).

• cp865 IBM 865 code page.

• applemac Macintosh encoding.

• next Next encoding.

• ansinew Windows 3.1 ANSI encoding, extension of Latin-1.

   • cp1252 Synonym for ansinew.

• cp1250 Windows 1250 (central and eastern Europe) code page.

Each encoding has an associated .def file, for example latin1.def which defines the behaviour of each input character, using the commands:

     \DeclareInputText{slot}{text}
     \DeclareInputMath{slot}{math}

This defines the input character slot to be the text material or math material respectively. For example, latin1.def defines slots "D6 (Æ) and "B5 (μ) by saying:

    \DeclareInputText{214}{\AE}
    \DeclareInputMath{181}{\mu}

Note that the commands should be robust, and should not be dependent on the output encoding. The same slot should not have both a text and a math declaration for it. (This restriction may be removed in future releases of inputenc).

The .def file may also define commands using the declarations:

\providecommand or \ProvideTextCommandDefault. For example:
    \ProvideTextCommandDefault{\textonequarter}{\ensuremath{\frac14}}
    \DeclareInputText{188}{\textonequarter}

The use of the ‘provide’ forms here will ensure that a better definition will not be over-written; their use is recommended since, in general, the best defintion depends on the fonts available.

See the documentation in fntguide.tex and ltoutenc.dtx for details of how to declare text commands.

2 Announcing the files

We announce the files:

 1 ⟨package⟩\NeedsTeXFormat{LaTeX2e}[1995/12/01]
 2 ⟨package⟩\message{beta test version}
 3 ⟨package⟩\ProvidesPackage{inputenc}
 4 ⟨ascii⟩ \ProvidesFile{ascii.def}
 5 ⟨latin1⟩ \ProvidesFile{latin1.def}
 6 ⟨latin2⟩ \ProvidesFile{latin2.def}
 7 ⟨latin3⟩ \ProvidesFile{latin3.def}
 8 ⟨latin5⟩ \ProvidesFile{latin5.def}
 9 ⟨decmulti⟩ \ProvidesFile{decmulti.def}
10 ⟨cp850⟩ \ProvidesFile{cp850.def}
11 ⟨cp852⟩ \ProvidesFile{cp852.def}

12 ⟨cp437⟩ \ProvidesFile{cp437.def}
13 ⟨cp437de⟩ \ProvidesFile{cp437de.def}
14 ⟨cp865⟩ \ProvidesFile{cp865.def}
15 ⟨applemac⟩ \ProvidesFile{applemac.def}
16 ⟨next⟩ \ProvidesFile{next.def}
17 ⟨ansinew⟩ \ProvidesFile{ansinew.def}
18 ⟨cp1252⟩ \ProvidesFile{cp1252.def}
19 ⟨cp1250⟩ \ProvidesFile{cp1250.def}
20 [1997/12/20 v0.95 Input encoding file
21  (test version: still liable to change)]

Now we insert a \makeatletter at the beginning of all .def files.

22 ⟨−package⟩\makeatletter

3 The package

Before we start with the code, an important comment is in order: as you may or may not know, the tabbing environment changes the definition of the commands \', \`, and \=. Outside such an environment these commands produce the corresponding accents, inside they are used for special text positioning, and the accents can be accessed using \a', \a`, and \a=. Therefore we must use the latter instead of the former in the second argument to \DeclareInputText, e.g.

(from latin1.def):
     \DeclareInputText{224}{\@tabacckludge`a}

The command \@tabacckludge is defined (in ltoutenc.dtx) in such a way that \@tabacckludge' will expand to the internal form of \'. Thus it is \' that is carried around internally (the same applies to the other two accent commands).

\DeclareInputText \DeclareInputMath \IeC These commands declare the expansion of an active character. The math declaration is the usual trick with \uppercase. The text declaration is sneakier, since in text space matters. We look to see if the definition ends in a macro, by checking whether it’s \meaning ends in a space. If it does, then we add an irrelevant \IeC and braces around the definition, in order to avoid any space after the active char being gobbled up once the text is written out to an auxiliary file.

The definition should contain only robust commands (and, for correct ligatures and kerning, they must be defined via the interfaces in the fontenc package).

23 ⟨∗package⟩

24 \def\DeclareInputMath#1{%
25 \@inpenc@test
26 \bgroup
27  \uccode`\~#1%
28  \uppercase{%
29 \egroup
30  \def~%
31 }%

32 }

33 \def\DeclareInputText#1#2{%
34 \def\@tempa{#2}%
35 \def\@tempb##1 ${}%
36 \edef\@tempa{\expandafter\@tempb\meaning\@tempa$ $}%
37 \ifx\@tempa\@empty
38  \DeclareInputMath{#1}{#2}%
39 \else
40  \DeclareInputMath{#1}{\IeC{#2}}%
41 \fi

42 }

The definition of \IeC was modified not to insert a \protect unless it is needed, this means it works in \hyphenation commands, and other such delicate places. It was then further changed to never insert a \protect as one is never needed; this makes it work in even more places.

This still needs some attention.

43 \DeclareRobustCommand{\IeC}[1]{#1}
44 \def\IeC{%
45 \ifx\protect\@typeset@protect
46  \expandafter\@firstofone
47 \else
48  \noexpand\IeC
49 \fi

50 }

\inputencoding This sets the encoding to be #1. It first sets all the characters 128–255 to be active (and sets their initial definition to be \@undefined). It now also does this for some ‘low’ codes below 32, but misses out Null, control-I, control-J, control-L and control-M.

It then inputs #1.def. This command should not be used in horizontal mode since space tokens within the file will produce unwanted spaces in the output.

But it first sets up a test that produces a warning message if no suitable definitions get read.

51 \def\inputencoding#1{%
52 \gdef\@inpenc@test{\global\let\@inpenc@test\relax}%
53 \ifvmode
54  \@inpenc@loop\^^A\^^H%
55  \@inpenc@loop\^^K\^^K%
56  \@inpenc@loop\^^N\^^_%
57  \@inpenc@loop\^^?\^^ff%
58  \input{#1.def}%
59 \fi
   If there have been no \DeclareInputText or \DeclareInputMath commands

read then something is amiss.

60 \ifx\@inpenc@test\relax\else
61  \PackageWarning{inputenc}%
62  {No characters defined\MessageBreak

63  by input encoding change to `#1'\MessageBreak}%
64 \fi

65 }

\@inpenc@loop Make characters #1 to #2 inclusive active and undefined.

66 \def\@inpenc@loop#1#2{%
67 \@tempcnta`#1\relax
68 \loop
69  \catcode\@tempcnta\active
70  \bgroup
71  \uccode`\~\@tempcnta
72  \uppercase{%
73  \egroup
74  \let~\@undefined
75  }%
76 \ifnum\@tempcnta<`#2\relax
77  \advance\@tempcnta\@ne
78 \repeat}

Then for each option, we input that encoding file.

79 \DeclareOption*{\inputencoding{\CurrentOption}}
80 \ProcessOptions

81 ⟨/package⟩

4 Default definitions for characters

Some input characters map to internal functions which are not in either the T1 or OT1 font encoding. For this reason default definitions are provided in the endoding file: these will be used unless some other output encoding is used which supports those glyphs. In some cases this default defintion has to be simply an error message.

Note that this works reasonably well only because the encoding files for both OT1 and T1 are loaded in the standard LaTeX format.

The name \textblacksquare is derived from the AMS symbol name since Adobe seem not to want this symbol. The default definition, as a rule, makes no claim to being a good design.

Some entries are repeated in case guards must all be on one line.

82 ⟨∗latin1 | decmulti | latin3 | latin5⟩

83 \ProvideTextCommandDefault{\textdegree}{\ensuremath{{^\circ}}}

84 ⟨/latin1 | decmulti | latin3 | latin5⟩ 85 ⟨∗cp850 | cp852 | cp865 | cp437 | cp437de | applemac | ansinew | cp1252 | cp1250 | next⟩

86 \ProvideTextCommandDefault{\textdegree}{\ensuremath{{^\circ}}}

87 ⟨/cp850 | cp852 | cp865 | cp437 | cp437de | applemac | ansinew | cp1252 | cp1250 | next⟩ 88 ⟨∗latin1 | decmulti | latin3 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩

89 \ProvideTextCommandDefault{\textonehalf}{\ensuremath{\frac12}}

90 ⟨/latin1 | decmulti | latin3 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩ 91 ⟨∗latin1 | decmulti | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩

 92 \ProvideTextCommandDefault{\textonequarter}{\ensuremath{\frac14}}

93 ⟨/latin1 | decmulti | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩ 94 ⟨∗latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩

 95 \ProvideTextCommandDefault{\textthreequarters}{\ensuremath{\frac34}}

96 ⟨/latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | next⟩ 97 ⟨∗applemac | cp850 | cp865 | cp437 | cp437de | ansinew | cp1252 | next⟩

 98 \ProvideTextCommandDefault{\textflorin}{\textit{f}}

99 ⟨/applemac | cp850 | cp865 | cp437 | cp437de | ansinew | cp1252 | next⟩ 100 ⟨∗cp865 | cp437 | cp437de⟩

101 \ProvideTextCommandDefault{\textpeseta}{Pt}

102 ⟨/cp865 | cp437 | cp437de⟩ 103 ⟨∗cp850 | cp852 | cp865 | cp437 | cp437de⟩

104 \ProvideTextCommandDefault{\textblacksquare}
105 {\vrule \@width .3em \@height .4em \@depth -.1em\relax}

106 ⟨/cp850 | cp852 | cp865 | cp437 | cp437de⟩

Some commands can’t be faked, so we have them generate an error message.

107 ⟨∗latin1 | decmulti | latin5 | cp850 | cp865 | cp437 | cp437de⟩

108 \ProvideTextCommandDefault{\textcent}
109 {\TextSymbolUnavailable\textcent}
110 \ProvideTextCommandDefault{\textyen}
111 {\TextSymbolUnavailable\textyen}

112 ⟨/latin1 | decmulti | latin5 | cp850 | cp865 | cp437 | cp437de⟩ 113 ⟨∗applemac | ansinew | cp1252⟩

114 \ProvideTextCommandDefault{\textcent}
115 {\TextSymbolUnavailable\textcent}
116 \ProvideTextCommandDefault{\textyen}
117 {\TextSymbolUnavailable\textyen}

118 ⟨/applemac | ansinew | cp1252⟩ 119 ⟨∗latin1 | decmulti | latin2 | latin3 | latin5 | cp850 | cp852 | cp865⟩

120 \ProvideTextCommandDefault{\textcurrency}
121 {\TextSymbolUnavailable\textcurrency}

122 ⟨/latin1 | decmulti | latin2 | latin3 | latin5 | cp850 | cp852 | cp865⟩ 123 ⟨∗applemac | ansinew | cp1252 | cp1250⟩

124 \ProvideTextCommandDefault{\textcurrency}
125 {\TextSymbolUnavailable\textcurrency}

126 ⟨/applemac | ansinew | cp1252 | cp1250⟩ 127 ⟨∗latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | cp1250⟩

128 \ProvideTextCommandDefault{\textbrokenbar}
129 {\TextSymbolUnavailable\textbrokenbar}

130 ⟨/latin1 | latin5 | cp850 | cp852 | ansinew | cp1252 | cp1250⟩ 131 ⟨∗latin3⟩

132 \ProvideTextCommandDefault{\textmalteseH}
133 {\TextSymbolUnavailable\textmalteseH}
134 \ProvideTextCommandDefault{\textmalteseh}
135 {\TextSymbolUnavailable\textmalteseh}

136 ⟨/latin3⟩ 137 ⟨∗next⟩

138 \ProvideTextCommandDefault{\textquotesingle}
139 {\TextSymbolUnavailable\textquotesingle}

140 ⟨/next⟩ 141 ⟨∗cp1250 | ansinew | cp1252 | applemac⟩

142 \ProvideTextCommandDefault{\textperthousand}
143 {\TextSymbolUnavailable\textperthousand}

144 ⟨/cp1250 | ansinew | cp1252 | applemac⟩

Characters that are supposed to be used only in math will be defined by \providecommand because L^{A}TEX 2ε assumes that the font encoding for math fonts is static.

145 ⟨∗latin1 | decmulti | latin5 | cp850 | ansinew | cp1252 | next⟩

146 \providecommand{\mathonesuperior}{{^1}}

147 ⟨/latin1 | decmulti | latin5 | cp850 | ansinew | cp1252 | next⟩ 148 ⟨∗latin1 | decmulti | latin3 | latin5 | cp850 | ansinew | cp1252 | next⟩

149 \providecommand{\maththreesuperior}{{^3}}

150 ⟨/latin1 | decmulti | latin3 | latin5 | cp850 | ansinew | cp1252 | next⟩ 151 ⟨∗latin1 | decmulti | latin3 | latin5 | cp850 | cp865 | cp437 | cp437de⟩

152 \providecommand{\mathtwosuperior}{{^2}}

153 ⟨/latin1 | decmulti | latin3 | latin5 | cp850 | cp865 | cp437 | cp437de⟩ 154 ⟨∗ansinew | cp1252 | next⟩

155 \providecommand{\mathtwosuperior}{{^2}}

156 ⟨/ansinew | cp1252 | next⟩ 157 ⟨∗cp865 | cp437 | cp437de⟩

158 \providecommand{\mathnsuperior}{{^n}}

159 ⟨/cp865 | cp437 | cp437de⟩ 160 ⟨∗latin1 | decmulti | latin5 | applemac | cp850 | cp852 | cp865 | cp437 | cp437de⟩

161 %|ansinew|cp1252|next>
162 \providecommand{\mathordmasculine}{{^o}}
163 \providecommand{\mathordfeminine}{{^a}}

164 ⟨/latin1 | decmulti | latin5 | applemac | cp850 | cp852 | cp865 | cp437 | cp437de⟩ 165 ⟨∗ansinew | cp1252 | next⟩

166 \providecommand{\mathordmasculine}{{^o}}
167 \providecommand{\mathordfeminine}{{^a}}

168 ⟨/ansinew | cp1252 | next⟩

5 The ASCII encoding

The ASCII encoding only allows characters in the range 32–127, so we only need to provide an empty .def file.

6 The ISO Latin-2 encoding

The ISO Latin-2 encoding file defines the characters in the ISO 8859-2 encoding. It was contributed by Petr Sojka (sojka@muni.cz) with small technical updates by Frank Mittelbach.

169 ⟨∗latin2⟩

170 \DeclareInputText{160}{\nobreakspace}
171 \DeclareInputText{176}{\textdegree}

172 \DeclareInputText{161}{\k A}
173 \DeclareInputText{177}{\k a}
174 \DeclareInputText{162}{\u{}}
175 \DeclareInputText{178}{\k\ }
176 \DeclareInputText{163}{\L}
177 \DeclareInputText{179}{\l}
178 \DeclareInputText{164}{\textcurrency}
179 \DeclareInputText{180}{\@tabacckludge'{}}
180 \DeclareInputText{165}{\v L}
181 \DeclareInputText{181}{\v l}
182 \DeclareInputText{166}{\@tabacckludge'S}
183 \DeclareInputText{182}{\@tabacckludge's}
184 \DeclareInputText{167}{\S}
185 \DeclareInputText{183}{\v{}}
186 \DeclareInputText{168}{\"{}}
187 \DeclareInputText{184}{\c\ }
188 \DeclareInputText{169}{\v S}
189 \DeclareInputText{185}{\v s}
190 \DeclareInputText{170}{\c S}
191 \DeclareInputText{186}{\c s}
192 \DeclareInputText{171}{\v T}
193 \DeclareInputText{187}{\v t}
194 \DeclareInputText{172}{\@tabacckludge'Z}
195 \DeclareInputText{188}{\@tabacckludge'z}
196 \DeclareInputText{173}{\-}
197 \DeclareInputText{189}{\H{}}
198 \DeclareInputText{174}{\v Z}
199 \DeclareInputText{190}{\v z}
200 \DeclareInputText{175}{\.Z}
201 \DeclareInputText{191}{\.z}
202 \DeclareInputText{192}{\@tabacckludge'R}
203 \DeclareInputText{208}{\DJ}
204 \DeclareInputText{193}{\@tabacckludge'A}
205 \DeclareInputText{209}{\@tabacckludge'N}
206 \DeclareInputText{194}{\^A}
207 \DeclareInputText{210}{\v N}
208 \DeclareInputText{195}{\u A}
209 \DeclareInputText{211}{\@tabacckludge'O}
210 \DeclareInputText{196}{\"A}
211 \DeclareInputText{212}{\^O}
212 \DeclareInputText{197}{\@tabacckludge'L}
213 \DeclareInputText{213}{\H O}
214 \DeclareInputText{198}{\@tabacckludge'C}
215 \DeclareInputText{214}{\"O}
216 \DeclareInputText{199}{\c C}
217 \DeclareInputMath{215}{\times}
218 \DeclareInputText{200}{\v C}
219 \DeclareInputText{216}{\v R}
220 \DeclareInputText{201}{\@tabacckludge'E}

221 \DeclareInputText{217}{\r U}
222 \DeclareInputText{202}{\k E}
223 \DeclareInputText{218}{\@tabacckludge'U}
224 \DeclareInputText{203}{\"E}
225 \DeclareInputText{219}{\H U}
226 \DeclareInputText{204}{\v E}
227 \DeclareInputText{220}{\"U}
228 \DeclareInputText{205}{\@tabacckludge'I}
229 \DeclareInputText{221}{\@tabacckludge'Y}
230 \DeclareInputText{206}{\^I}
231 \DeclareInputText{222}{\c T}
232 \DeclareInputText{207}{\v D}
233 \DeclareInputText{223}{\ss}
234 \DeclareInputText{224}{\@tabacckludge'r}
235 \DeclareInputText{240}{\dj}
236 \DeclareInputText{225}{\@tabacckludge'a}
237 \DeclareInputText{241}{\@tabacckludge'n}
238 \DeclareInputText{226}{\^a}
239 \DeclareInputText{242}{\v n}
240 \DeclareInputText{227}{\u a}
241 \DeclareInputText{243}{\@tabacckludge'o}
242 \DeclareInputText{228}{\"a}
243 \DeclareInputText{244}{\^o}
244 \DeclareInputText{229}{\@tabacckludge'l}
245 \DeclareInputText{245}{\H o}
246 \DeclareInputText{230}{\@tabacckludge'c}
247 \DeclareInputText{246}{\"o}
248 \DeclareInputText{231}{\c c}
249 \DeclareInputMath{247}{\div}
250 \DeclareInputText{232}{\v c}
251 \DeclareInputText{248}{\v r}
252 \DeclareInputText{233}{\@tabacckludge'e}
253 \DeclareInputText{249}{\r u}
254 \DeclareInputText{234}{\k e}
255 \DeclareInputText{250}{\@tabacckludge'u}
256 \DeclareInputText{235}{\"e}
257 \DeclareInputText{251}{\H u}
258 \DeclareInputText{236}{\v e}
259 \DeclareInputText{252}{\"u}
260 \DeclareInputText{237}{\@tabacckludge'\i}
261 \DeclareInputText{253}{\@tabacckludge'y}
262 \DeclareInputText{238}{\^\i}
263 \DeclareInputText{254}{\c t}
264 \DeclareInputText{239}{\v d}
265 \DeclareInputText{255}{\.{}}

266 ⟨/latin2⟩

7 The ISO Latin-3 encoding

The ISO Latin-3 encoding file defines the characters in the ISO 8859-3 encoding. It was contributed by by Jörg Knappen (joerg.knappen@uni-mainz.de) and adapted by Chris Rowley.

It can be used for general purpose applications in typical office environments in the following languages: Afrikaans, Catalan, English, Esperanto, French, Gallician, German, Italian, Maltese, and Turkish.

267 ⟨∗latin3⟩

268 \DeclareInputText{160}{\nobreakspace}
269 \DeclareInputText{176}{\textdegree}
270 \DeclareInputText{161}{\textmalteseH}
271 \DeclareInputText{177}{\textmalteseh}
272 \DeclareInputText{162}{\u{}}
273 \DeclareInputMath{178}{\mathtwosuperior}
274 \DeclareInputText{163}{\pounds}
275 \DeclareInputMath{179}{\maththreesuperior}
276 \DeclareInputText{164}{\textcurrency}
277 \DeclareInputText{180}{\@tabacckludge'{}}
278 % \DeclareInputText{165}{\notdef}
279 \DeclareInputMath{181}{\mu}
280 \DeclareInputText{166}{\^H}
281 % NOT: \DeclareInputText{182}{h\llap{\^{}}} % \^h would be too tall
282 \DeclareInputText{182}{\^h}
283 \DeclareInputText{167}{\S}
284 \DeclareInputText{183}{\textperiodcentered}
285 \DeclareInputText{168}{\"{}}
286 \DeclareInputText{184}{\c\ }
287 \DeclareInputText{169}{\.I}
288 \DeclareInputText{185}{\i}
289 \DeclareInputText{170}{\c S}
290 \DeclareInputText{186}{\c s}
291 \DeclareInputText{171}{\u G}
292 \DeclareInputText{187}{\u g}
293 \DeclareInputText{172}{\^J}
294 \DeclareInputText{188}{\^\j}
295 \DeclareInputText{173}{\-}
296 \DeclareInputText{189}{\textonehalf}
297 % \DeclareInputText{174}{\notdef}
298 % \DeclareInputText{190}{\notdef}
299 \DeclareInputText{175}{\.Z}
300 \DeclareInputText{191}{\.z}
301 \DeclareInputText{192}{\@tabacckludge`A}
302 % \DeclareInputText{208}{\notdef}
303 \DeclareInputText{193}{\@tabacckludge'A}
304 \DeclareInputText{209}{\~N}
305 \DeclareInputText{194}{\^A}
306 \DeclareInputText{210}{\@tabacckludge`O}
307 % \DeclareInputText{195}{\notdef}

308 \DeclareInputText{211}{\@tabacckludge'O}
309 \DeclareInputText{196}{\"A}
310 \DeclareInputText{212}{\^O}
311 \DeclareInputText{197}{\.C}
312 \DeclareInputText{213}{\.G}
313 \DeclareInputText{198}{\^C}
314 \DeclareInputText{214}{\"O}
315 \DeclareInputText{199}{\c C}
316 \DeclareInputMath{215}{\times}
317 \DeclareInputText{200}{\@tabacckludge`E}
318 \DeclareInputText{216}{\^G}
319 \DeclareInputText{201}{\@tabacckludge'E}
320 \DeclareInputText{217}{\@tabacckludge`U}
321 \DeclareInputText{202}{\^E}
322 \DeclareInputText{218}{\@tabacckludge'U}
323 \DeclareInputText{203}{\"E}
324 \DeclareInputText{219}{\^U}
325 \DeclareInputText{204}{\@tabacckludge`I}
326 \DeclareInputText{220}{\"U}
327 \DeclareInputText{205}{\@tabacckludge'I}
328 \DeclareInputText{221}{\u U}
329 \DeclareInputText{206}{\^I}
330 \DeclareInputText{222}{\^S}
331 \DeclareInputText{207}{\"I}
332 \DeclareInputText{223}{\ss}
333 \DeclareInputText{224}{\@tabacckludge`a}
334 % \DeclareInputText{240}{\notdef}
335 \DeclareInputText{225}{\@tabacckludge'a}
336 \DeclareInputText{241}{\~n}
337 \DeclareInputText{226}{\^a}
338 \DeclareInputText{242}{\@tabacckludge`o}
339 % \DeclareInputText{227}{\notdef}
340 \DeclareInputText{243}{\@tabacckludge'o}
341 \DeclareInputText{228}{\"a}
342 \DeclareInputText{244}{\^o}
343 \DeclareInputText{229}{\.c}
344 \DeclareInputText{245}{\.g}
345 \DeclareInputText{230}{\^c}
346 \DeclareInputText{246}{\"o}
347 \DeclareInputText{231}{\c c}
348 \DeclareInputMath{247}{\div}
349 \DeclareInputText{232}{\@tabacckludge`e}
350 \DeclareInputText{248}{\^g}
351 \DeclareInputText{233}{\@tabacckludge'e}
352 \DeclareInputText{249}{\@tabacckludge`u}
353 \DeclareInputText{234}{\^e}
354 \DeclareInputText{250}{\@tabacckludge'u}
355 \DeclareInputText{235}{\"e}
356 \DeclareInputText{251}{\^u}

357 \DeclareInputText{236}{\@tabacckludge`\i}
358 \DeclareInputText{252}{\"u}
359 \DeclareInputText{237}{\@tabacckludge'\i}
360 \DeclareInputText{253}{\u u}
361 \DeclareInputText{238}{\^\i}
362 \DeclareInputText{254}{\^s}
363 \DeclareInputText{239}{\"\i}
364 \DeclareInputText{255}{\.{}}

365 ⟨/latin3⟩

8 The ISO Latin-5 encoding

The ISO Latin-5 encoding file defines the characters in the ISO 8859-5 encoding. It was contributed by H. Turgut Uyar (uyar@cs.itu.edu.tr); it is used for Turkish.

366 ⟨∗latin5⟩

367 \DeclareInputText{160}{\nobreakspace}
368 \DeclareInputText{176}{\textdegree}
369 \DeclareInputText{161}{\textexclamdown}
370 \DeclareInputMath{177}{\pm}
371 \DeclareInputText{162}{\textcent}
372 \DeclareInputMath{178}{\mathtwosuperior}
373 \DeclareInputText{163}{\pounds}
374 \DeclareInputMath{179}{\maththreesuperior}
375 \DeclareInputText{164}{\textcurrency}
376 \DeclareInputText{180}{\@tabacckludge'{}}
377 \DeclareInputText{165}{\textyen}
378 \DeclareInputMath{181}{\mu}
379 \DeclareInputText{166}{\textbrokenbar}
380 \DeclareInputText{182}{\P}
381 \DeclareInputText{167}{\S}
382 \DeclareInputText{183}{\textperiodcentered}
383 \DeclareInputText{168}{\"{}}
384 \DeclareInputText{184}{\c\ }
385 \DeclareInputText{169}{\copyright}
386 \DeclareInputMath{185}{\mathonesuperior}
387 \DeclareInputMath{170}{\mathordfeminine}
388 \DeclareInputMath{186}{\mathordmasculine}
389 \DeclareInputText{171}{\guillemotleft}
390 \DeclareInputText{187}{\guillemotright}
391 \DeclareInputMath{172}{\lnot}
392 \DeclareInputText{188}{\textonequarter}
393 \DeclareInputText{173}{\-}
394 \DeclareInputText{189}{\textonehalf}
395 \DeclareInputText{174}{\textregistered}
396 \DeclareInputText{190}{\textthreequarters}
397 \DeclareInputText{175}{\@tabacckludge={}}
398 \DeclareInputText{191}{\textquestiondown}
399 \DeclareInputText{192}{\@tabacckludge`A}

400 \DeclareInputText{208}{\u G}
401 \DeclareInputText{193}{\@tabacckludge'A}
402 \DeclareInputText{209}{\~N}
403 \DeclareInputText{194}{\^A}
404 \DeclareInputText{210}{\@tabacckludge`O}
405 \DeclareInputText{195}{\~A}
406 \DeclareInputText{211}{\@tabacckludge'O}
407 \DeclareInputText{196}{\"A}
408 \DeclareInputText{212}{\^O}
409 \DeclareInputText{197}{\r A}
410 \DeclareInputText{213}{\~O}
411 \DeclareInputText{198}{\AE}
412 \DeclareInputText{214}{\"O}
413 \DeclareInputText{199}{\c C}
414 \DeclareInputMath{215}{\times}
415 \DeclareInputText{200}{\@tabacckludge`E}
416 \DeclareInputText{216}{\O}
417 \DeclareInputText{201}{\@tabacckludge'E}
418 \DeclareInputText{217}{\@tabacckludge`U}
419 \DeclareInputText{202}{\^E}
420 \DeclareInputText{218}{\@tabacckludge'U}
421 \DeclareInputText{203}{\"E}
422 \DeclareInputText{219}{\^U}
423 \DeclareInputText{204}{\@tabacckludge`I}
424 \DeclareInputText{220}{\"U}
425 \DeclareInputText{205}{\@tabacckludge'I}
426 \DeclareInputText{221}{\.I}
427 \DeclareInputText{206}{\^I}
428 \DeclareInputText{222}{\c S}
429 \DeclareInputText{207}{\"I}
430 \DeclareInputText{223}{\ss}
431 \DeclareInputText{224}{\@tabacckludge`a}
432 \DeclareInputText{240}{\u g}
433 \DeclareInputText{225}{\@tabacckludge'a}
434 \DeclareInputText{241}{\~n}
435 \DeclareInputText{226}{\^a}
436 \DeclareInputText{242}{\@tabacckludge`o}
437 \DeclareInputText{227}{\~a}
438 \DeclareInputText{243}{\@tabacckludge'o}
439 \DeclareInputText{228}{\"a}
440 \DeclareInputText{244}{\^o}
441 \DeclareInputText{229}{\r a}
442 \DeclareInputText{245}{\~o}
443 \DeclareInputText{230}{\ae}
444 \DeclareInputText{246}{\"o}
445 \DeclareInputText{231}{\c c}
446 \DeclareInputMath{247}{\div}
447 \DeclareInputText{232}{\@tabacckludge`e}
448 \DeclareInputText{248}{\o}

449 \DeclareInputText{233}{\@tabacckludge'e}
450 \DeclareInputText{249}{\@tabacckludge`u}
451 \DeclareInputText{234}{\^e}
452 \DeclareInputText{250}{\@tabacckludge'u}
453 \DeclareInputText{235}{\"e}
454 \DeclareInputText{251}{\^u}
455 \DeclareInputText{236}{\@tabacckludge`\i}
456 \DeclareInputText{252}{\"u}
457 \DeclareInputText{237}{\@tabacckludge'\i}
458 \DeclareInputText{253}{\i }
459 \DeclareInputText{238}{\^\i}
460 \DeclareInputText{254}{\c s}
461 \DeclareInputText{239}{\"\i}
462 \DeclareInputText{255}{\"y}

463 ⟨/latin5⟩

9 DEC Multinational Character Set

The DECMultinational character set, used by the OpenVMS operating system, is slightly different from the ISO Latin 1 character set.

Reference: Digital Equipment Corporation VT330/VT340 Programmer Reference Manual, Volume 1: Text Programming, page 22.

This encoding was provided by M.Y. Chartoire IPNL-IN2P3

 m.chartoire@ipnl.in2p3.fr

464 ⟨∗decmulti⟩

465 \DeclareInputText{176}{\textdegree}
466 \DeclareInputText{161}{\textexclamdown}
467 \DeclareInputMath{177}{\pm}
468 \DeclareInputText{162}{\textcent}
469 \DeclareInputMath{178}{\mathtwosuperior}
470 \DeclareInputText{163}{\pounds}
471 \DeclareInputMath{179}{\maththreesuperior}
472 \DeclareInputText{165}{\textyen}
473 \DeclareInputMath{181}{\mu}
474 \DeclareInputText{182}{\P}
475 \DeclareInputText{167}{\S}
476 \DeclareInputText{183}{\textperiodcentered}
477 \DeclareInputText{168}{\textcurrency}
478 \DeclareInputText{169}{\copyright}
479 \DeclareInputMath{185}{\mathonesuperior}
480 \DeclareInputMath{170}{\mathordfeminine}
481 \DeclareInputMath{186}{\mathordmasculine}
482 \DeclareInputText{171}{\guillemotleft}
483 \DeclareInputText{187}{\guillemotright}
484 \DeclareInputText{188}{\textonequarter}
485 \DeclareInputText{189}{\textonehalf}
486 \DeclareInputText{191}{\textquestiondown}

487 \DeclareInputText{192}{\@tabacckludge`A}
488 \DeclareInputText{193}{\@tabacckludge'A}
489 \DeclareInputText{209}{\~N}
490 \DeclareInputText{194}{\^A}
491 \DeclareInputText{210}{\@tabacckludge`O}
492 \DeclareInputText{195}{\~A}
493 \DeclareInputText{211}{\@tabacckludge'O}
494 \DeclareInputText{196}{\"A}
495 \DeclareInputText{212}{\^O}
496 \DeclareInputText{197}{\r A}
497 \DeclareInputText{213}{\~O}
498 \DeclareInputText{198}{\AE}
499 \DeclareInputText{214}{\"O}
500 \DeclareInputText{199}{\c C}
501 \DeclareInputText{215}{\OE}
502 \DeclareInputText{200}{\@tabacckludge`E}
503 \DeclareInputText{216}{\O}
504 \DeclareInputText{201}{\@tabacckludge'E}
505 \DeclareInputText{217}{\@tabacckludge`U}
506 \DeclareInputText{202}{\^E}
507 \DeclareInputText{218}{\@tabacckludge'U}
508 \DeclareInputText{203}{\"E}
509 \DeclareInputText{219}{\^U}
510 \DeclareInputText{204}{\@tabacckludge`I}
511 \DeclareInputText{220}{\"U}
512 \DeclareInputText{205}{\@tabacckludge'I}
513 \DeclareInputText{221}{\"Y}
514 \DeclareInputText{206}{\^I}
515 \DeclareInputText{207}{\"I}
516 \DeclareInputText{223}{\ss}
517 \DeclareInputText{224}{\@tabacckludge`a}
518 \DeclareInputText{225}{\@tabacckludge'a}
519 \DeclareInputText{241}{\~n}
520 \DeclareInputText{226}{\^a}
521 \DeclareInputText{242}{\@tabacckludge`o}
522 \DeclareInputText{227}{\~a}
523 \DeclareInputText{243}{\@tabacckludge'o}
524 \DeclareInputText{228}{\"a}
525 \DeclareInputText{244}{\^o}
526 \DeclareInputText{229}{\r a}
527 \DeclareInputText{245}{\~o}
528 \DeclareInputText{230}{\ae}
529 \DeclareInputText{246}{\"o}
530 \DeclareInputText{231}{\c c}
531 \DeclareInputText{247}{\oe}
532 \DeclareInputText{232}{\@tabacckludge`e}
533 \DeclareInputText{248}{\o}
534 \DeclareInputText{233}{\@tabacckludge'e}
535 \DeclareInputText{249}{\@tabacckludge`u}

536 \DeclareInputText{234}{\^e}
537 \DeclareInputText{250}{\@tabacckludge'u}
538 \DeclareInputText{235}{\"e}
539 \DeclareInputText{251}{\^u}
540 \DeclareInputText{236}{\@tabacckludge`\i}
541 \DeclareInputText{252}{\"u}
542 \DeclareInputText{237}{\@tabacckludge'\i}
543 \DeclareInputText{253}{\"y}
544 \DeclareInputText{238}{\^\i}
545 \DeclareInputText{239}{\"\i}

546 ⟨/decmulti⟩

10 The IBM code page 850

This input encoding was contributed by Timo Knuutila (knuutila@cs.utu.fi), and edited by Christian Bartels (ii140ba@vm1.rz.rwth-aachen.de).

The DOS graphics ‘letters’ and a few other positions are ignored (left undefined).

547 ⟨∗cp850⟩

548 \DeclareInputText{128}{\c C}
549 \DeclareInputText{144}{\@tabacckludge'E}
550 \DeclareInputText{129}{\"u}
551 \DeclareInputText{145}{\ae}
552 \DeclareInputText{130}{\@tabacckludge'e}
553 \DeclareInputText{146}{\AE}
554 \DeclareInputText{131}{\^a}
555 \DeclareInputText{147}{\^o}
556 \DeclareInputText{132}{\"a}
557 \DeclareInputText{148}{\"o}
558 \DeclareInputText{133}{\@tabacckludge`a}
559 \DeclareInputText{149}{\@tabacckludge`o}
560 \DeclareInputText{134}{\r a}
561 \DeclareInputText{150}{\^u}
562 \DeclareInputText{135}{\c c}
563 \DeclareInputText{151}{\@tabacckludge`u}
564 \DeclareInputText{136}{\^e}
565 \DeclareInputText{152}{\"y}
566 \DeclareInputText{137}{\"e}
567 \DeclareInputText{153}{\"O}
568 \DeclareInputText{138}{\@tabacckludge`e}
569 \DeclareInputText{154}{\"U}
570 \DeclareInputText{139}{\"\i}
571 \DeclareInputText{155}{\o}
572 \DeclareInputText{140}{\^\i}
573 \DeclareInputText{156}{\pounds}
574 \DeclareInputText{141}{\@tabacckludge`\i}
575 \DeclareInputText{157}{\O}
576 \DeclareInputText{142}{\"A}

577 \DeclareInputMath{158}{\times}
578 \DeclareInputText{143}{\r A}
579 \DeclareInputText{159}{\textflorin}
580 \DeclareInputText{160}{\@tabacckludge'a}
581 \DeclareInputText{161}{\@tabacckludge'\i}
582 \DeclareInputText{162}{\@tabacckludge'o}
583 \DeclareInputText{163}{\@tabacckludge'u}
584 \DeclareInputText{164}{\~n}               %% "B0-"B4: DG
585 \DeclareInputText{165}{\~N}
586 \DeclareInputText{181}{\@tabacckludge'A}
587 \DeclareInputMath{166}{\mathordfeminine}
588 \DeclareInputText{182}{\^A}
589 \DeclareInputMath{167}{\mathordmasculine}
590 \DeclareInputText{183}{\@tabacckludge`A}
591 \DeclareInputText{168}{\textquestiondown}
592 \DeclareInputText{184}{\copyright}
593 \DeclareInputText{169}{\textregistered}
594 \DeclareInputMath{170}{\lnot}
595 \DeclareInputText{171}{\textonehalf}
596 \DeclareInputText{172}{\textonequarter}   %% "B9-"BC: DG
597 \DeclareInputText{173}{\textexclamdown}
598 \DeclareInputText{189}{\textcent}
599 \DeclareInputText{174}{\guillemotleft}
600 \DeclareInputText{190}{\textyen}
601 \DeclareInputText{175}{\guillemotright}
602 % "BF-"C5: DG
603 \DeclareInputText{208}{\dh}
604 \DeclareInputText{209}{\DH}
605 \DeclareInputText{210}{\^E}
606 \DeclareInputText{211}{\"E}
607 \DeclareInputText{212}{\@tabacckludge`E}
608 \DeclareInputText{213}{\i}
609 \DeclareInputText{198}{\~a}
610 \DeclareInputText{214}{\@tabacckludge'I}
611 \DeclareInputText{199}{\~A}
612 \DeclareInputText{215}{\^I}
613 \DeclareInputText{216}{\"I}
614 % "C8-"CE: DG
615 % "D9-"DC: DG
616 \DeclareInputText{221}{\textbrokenbar}
617 \DeclareInputText{222}{\@tabacckludge`I}
618 \DeclareInputText{207}{\textcurrency}  %% "DF: DG
619 \DeclareInputText{224}{\@tabacckludge'O}
620 \DeclareInputText{240}{\textendash}
621 \DeclareInputText{225}{\ss}
622 \DeclareInputMath{241}{\pm}
623 \DeclareInputText{226}{\^O}
624 % "F2: DG (not double underline, or equals?)
625 \DeclareInputText{227}{\@tabacckludge`O}

626 \DeclareInputText{243}{\textthreequarters}
627 \DeclareInputText{228}{\~o}
628 \DeclareInputText{244}{\P}
629 \DeclareInputText{229}{\~O}
630 \DeclareInputText{245}{\S}
631 \DeclareInputMath{230}{\mu}
632 \DeclareInputMath{246}{\div}
633 \DeclareInputText{231}{\th}
634 \DeclareInputText{247}{\c\ }
635 \DeclareInputText{232}{\TH}
636 \DeclareInputText{248}{\textdegree}
637 \DeclareInputText{233}{\@tabacckludge'U}
638 \DeclareInputText{249}{\"{}}
639 \DeclareInputText{234}{\^U}
640 \DeclareInputText{250}{\textperiodcentered}
641 \DeclareInputText{235}{\@tabacckludge`U}
642 \DeclareInputMath{251}{\mathonesuperior}
643 \DeclareInputText{236}{\@tabacckludge'y}
644 \DeclareInputMath{252}{\maththreesuperior}
645 \DeclareInputText{237}{\@tabacckludge'Y}
646 \DeclareInputMath{253}{\mathtwosuperior}
647 \DeclareInputText{238}{\@tabacckludge={}}
648 \DeclareInputText{254}{\textblacksquare} % right name?
649 \DeclareInputText{239}{\@tabacckludge'{}}
650 \DeclareInputText{255}{\nobreakspace}

651 ⟨/cp850⟩

11 The IBM code page 852

This input encoding was contributed by Petr Sojka (sojka@Muni.cz).

652 ⟨∗cp852⟩

653 \DeclareInputText{128}{\c C}
654 \DeclareInputText{144}{\@tabacckludge'E}
655 \DeclareInputText{129}{\"u}
656 \DeclareInputText{145}{\@tabacckludge'L}
657 \DeclareInputText{130}{\@tabacckludge'e}
658 \DeclareInputText{146}{\@tabacckludge'l}
659 \DeclareInputText{131}{\^a}
660 \DeclareInputText{147}{\^o}
661 \DeclareInputText{132}{\"a}
662 \DeclareInputText{148}{\"o}
663 \DeclareInputText{133}{\r u}
664 \DeclareInputText{149}{\v L}
665 \DeclareInputText{134}{\@tabacckludge'c}
666 \DeclareInputText{150}{\v l}
667 \DeclareInputText{135}{\c c}
668 \DeclareInputText{151}{\@tabacckludge'S}
669 \DeclareInputText{136}{\l}

670 \DeclareInputText{152}{\@tabacckludge's}
671 \DeclareInputText{137}{\"e}
672 \DeclareInputText{153}{\"O}
673 \DeclareInputText{138}{\H O}
674 \DeclareInputText{154}{\"U}
675 \DeclareInputText{139}{\H o}
676 \DeclareInputText{155}{\v T}
677 \DeclareInputText{140}{\^\i}
678 \DeclareInputText{156}{\v t}
679 \DeclareInputText{141}{\@tabacckludge'Z}
680 \DeclareInputText{157}{\L}
681 \DeclareInputText{142}{\"A}
682 \DeclareInputMath{158}{\times}
683 \DeclareInputText{143}{\@tabacckludge'C}
684 \DeclareInputText{159}{\v c}
685 \DeclareInputText{160}{\@tabacckludge'a}
686 \DeclareInputText{161}{\@tabacckludge'\i}
687 \DeclareInputText{162}{\@tabacckludge'o}
688 \DeclareInputText{163}{\@tabacckludge'u}
689 \DeclareInputText{164}{\k A}
690 \DeclareInputText{165}{\k a}
691 \DeclareInputText{166}{\v Z}
692 \DeclareInputText{167}{\v z}
693 \DeclareInputText{168}{\k E}
694 \DeclareInputText{169}{\k e}
695 \DeclareInputText{170}{\nobreakspace}  % or empty??
696 \DeclareInputText{171}{\@tabacckludge'z}
697 \DeclareInputText{172}{\v C}
698 \DeclareInputText{173}{\c s}
699 \DeclareInputText{174}{\guillemotleft}
700 \DeclareInputText{175}{\guillemotright}
701 \DeclareInputText{181}{\@tabacckludge'A}
702 \DeclareInputText{182}{\^A}
703 \DeclareInputText{183}{\v E}
704 \DeclareInputText{184}{\c S}
705 \DeclareInputText{189}{\.Z}
706 \DeclareInputText{190}{\.z}
707 \DeclareInputText{198}{\u A}
708 \DeclareInputText{199}{\u a}
709 \DeclareInputText{207}{\textcurrency}
710 \DeclareInputText{208}{\dj}
711 \DeclareInputText{209}{\DJ}
712 \DeclareInputText{210}{\v D}
713 \DeclareInputText{211}{\"E}
714 \DeclareInputText{212}{\@tabacckludge'd}
715 \DeclareInputText{213}{\v N}
716 \DeclareInputText{214}{\@tabacckludge'I}
717 \DeclareInputText{215}{\^I}
718 \DeclareInputText{216}{\v e}

719 \DeclareInputText{221}{\c T}
720 \DeclareInputText{222}{\r U}
721 \DeclareInputText{224}{\@tabacckludge'O}
722 \DeclareInputText{240}{\textendash}
723 \DeclareInputText{225}{\ss}
724 \DeclareInputText{241}{\H{}}
725 \DeclareInputText{226}{\^O}
726 \DeclareInputText{242}{\k\ }
727 \DeclareInputText{227}{\@tabacckludge'N}
728 \DeclareInputText{243}{\v{}}
729 \DeclareInputText{228}{\@tabacckludge'n}
730 \DeclareInputText{244}{\u{}}
731 \DeclareInputText{229}{\v n}
732 \DeclareInputText{245}{\S}
733 \DeclareInputText{230}{\v S}
734 \DeclareInputMath{246}{\div}
735 \DeclareInputText{231}{\v s}
736 \DeclareInputText{247}{\c\ }
737 \DeclareInputText{232}{\@tabacckludge'R}
738 \DeclareInputText{248}{\textdegree}
739 \DeclareInputText{233}{\@tabacckludge'U}
740 \DeclareInputText{249}{\"{}}
741 \DeclareInputText{234}{\@tabacckludge'r}
742 \DeclareInputText{250}{\.{}}
743 \DeclareInputText{235}{\H U}
744 \DeclareInputText{251}{\H u}
745 \DeclareInputText{236}{\@tabacckludge'y}
746 \DeclareInputText{252}{\v R}
747 \DeclareInputText{237}{\@tabacckludge'Y}
748 \DeclareInputText{253}{\v r}
749 \DeclareInputText{238}{\c t}
750 \DeclareInputText{254}{\textblacksquare} % right name?
751 \DeclareInputText{239}{\@tabacckludge'{}}
752 \DeclareInputText{255}{\nobreakspace}

753 ⟨/cp852⟩

12 The IBM codepages 437 and 865

This input encoding is based on work by

 Volker Kunert (volker@numsun1.mathematik.uni-halle.de)
 and bontus@al6000.physik.uni-siegen.de.

The changes for cp865 are based on work by Søren Sandmann (sandmann@ daimi.aau.dk), with thanks to them all.

The DOS graphics ‘letters’ and a few other positions are ignored (left undefined).

Unfortunately, in cp437 there is no agreement as to whether slot E1 should be ‘β’ or ‘ß’, so we provide two variants, one (cp437) with ‘β’ and one (cp437de) with ‘ß’.

754 ⟨∗cp437 | cp437de | cp865⟩

755 \DeclareInputText{128}{{\c C}}
756 \DeclareInputText{129}{\"u}
757 \DeclareInputText{130}{\@tabacckludge'e}
758 \DeclareInputText{131}{\^a}
759 \DeclareInputText{132}{\"a}
760 \DeclareInputText{133}{\@tabacckludge`a}
761 \DeclareInputText{134}{\r a}
762 \DeclareInputText{135}{{\c c}}
763 \DeclareInputText{136}{\^e}
764 \DeclareInputText{137}{\"e}
765 \DeclareInputText{138}{\@tabacckludge`e}
766 \DeclareInputText{139}{\"\i}
767 \DeclareInputText{140}{\^\i}
768 \DeclareInputText{141}{\@tabacckludge`\i}
769 \DeclareInputText{142}{\"A}
770 \DeclareInputText{143}{\r A}
771 \DeclareInputText{144}{\@tabacckludge'E}
772 \DeclareInputText{145}{\ae}
773 \DeclareInputText{146}{\AE}
774 \DeclareInputText{147}{\^o}
775 \DeclareInputText{148}{\"o}
776 \DeclareInputText{149}{\@tabacckludge`o}
777 \DeclareInputText{150}{\^u}
778 \DeclareInputText{151}{\@tabacckludge`u}
779 \DeclareInputText{152}{\"y}
780 \DeclareInputText{153}{\"O}
781 \DeclareInputText{154}{\"U}

782 ⟨/cp437 | cp437de | cp865⟩ 783 ⟨∗cp437 | cp437de⟩

784 \DeclareInputText{155}{\textcent}
785 \DeclareInputText{156}{\pounds}
786 \DeclareInputText{157}{\textyen}

787 ⟨/cp437 | cp437de⟩ 788 ⟨∗cp865⟩

789 \DeclareInputText{155}{\o}
790 \DeclareInputText{156}{\pounds}
791 \DeclareInputText{157}{\O}

792 ⟨/cp865⟩ 793 ⟨∗cp437 | cp437de | cp865⟩

794 \DeclareInputText{158}{\textpeseta} % Pt
795 \DeclareInputText{159}{\textflorin}
796 \DeclareInputText{160}{\@tabacckludge'a} % 160
797 \DeclareInputText{161}{\@tabacckludge'\i}
798 \DeclareInputText{162}{\@tabacckludge'o}
799 \DeclareInputText{163}{\@tabacckludge'u}
800 \DeclareInputText{164}{\~n}
801 \DeclareInputText{165}{\~N}

802 \DeclareInputMath{166}{\mathordfeminine}
803 \DeclareInputMath{167}{\mathordmasculine}
804 \DeclareInputText{168}{\textquestiondown}
805 %\DeclareInputText{169}{} % left upper corner
806 \DeclareInputMath{170}{\lnot}
807 \DeclareInputText{171}{\textonehalf}
808 \DeclareInputText{172}{\textonequarter}
809 \DeclareInputText{173}{\textexclamdown}
810 \DeclareInputText{174}{\guillemotleft}
811 ⟨cp437 | cp437de⟩\DeclareInputText{175}{\guillemotright}
812 ⟨cp865⟩\DeclareInputText{175}{\textcurrency}
813 %\DeclareInputText{176}{\textlightgraybox}
814 %\DeclareInputText{177}{\textgraybox}
815 %\DeclareInputText{178}{\textdarkgraybox}
816 %\DeclareInputMath{179}{} % vertical bar
817 %\DeclareInputText{180}{} % vertical bar with branch to left
818 %\DeclareInputText{181}{} % vertical bar with double branch to left
819 %\DeclareInputText{182}{} % double bar with single branch to left
820 %\DeclareInputText{183}{} % graphic
821 %\DeclareInputText{184}{} % graphic
822 %\DeclareInputMath{185}{} % vertical double bar with branch to left
823 %\DeclareInputMath{186}{} % vertical double bar
824 %\DeclareInputText{187}{} % double upper right corner
825 %\DeclareInputMath{188}{} % double lower right corner
826 %\DeclareInputText{189}{} % graphic
827 %\DeclareInputText{190}{} % graphic
828 %\DeclareInputMath{191}{\ensuremath{\rceil}}
829 %\DeclareInputMath{192}{\ensuremath{\lfloor}}
830 %\DeclareInputText{193}{} % dash with branch up
831 %\DeclareInputText{194}{} % dash with branch down
832 %\DeclareInputText{195}{} % vertical bar with branch to right
833 %\DeclareInputText{196}{} % horizontal bar, not endash
834 %\DeclareInputText{197}{} % vertical bar crossed with dash
835 %\DeclareInputText{198}{} % graphic
836 %\DeclareInputText{199}{} % graphic
837 %\DeclareInputText{200}{} % double lower left corner
838 %\DeclareInputText{201}{} % double upper left corner
839 %\DeclareInputText{202}{} % double dash with branch up
840 %\DeclareInputText{203}{} % double dash with branch down
841 %\DeclareInputText{204}{} % double bar with branch right
842 %\DeclareInputText{205}{=} % double dash
843 %\DeclareInputText{206}{} % double bar crossing double dash
844 %\DeclareInputText{207}{} % graphic
845 %\DeclareInputMath{208}{}
846 %\DeclareInputText{209}{}
847 %\DeclareInputText{210}{}
848 %\DeclareInputText{211}{}
849 %\DeclareInputText{212}{}
850 %\DeclareInputText{213}{}

851 %\DeclareInputText{214}{}
852 %\DeclareInputText{215}{}
853 %\DeclareInputText{216}{}
854 %\DeclareInputMath{217}{\ensuremath{\rfloor}} % lower right corner
855 %\DeclareInputMath{218}{\ensuremath{\lceil}}  % upper left corner
856 %\DeclareInputText{219}{} % black box
857 %\DeclareInputText{220}{} % lower half of black box
858 %\DeclareInputText{221}{} % left bar
859 %\DeclareInputText{222}{} % right bar
860 %\DeclareInputText{223}{} % upper half of black box
861 \DeclareInputMath{224}{\alpha}

862 ⟨/cp437 | cp437de | cp865⟩

863 ⟨cp437 | cp865⟩\DeclareInputMath{225}{\beta}
864 ⟨cp437de⟩\DeclareInputText{225}{\ss}

865 ⟨∗cp437 | cp437de | cp865⟩

866 \DeclareInputMath{226}{\Gamma}
867 \DeclareInputMath{227}{\pi}
868 \DeclareInputMath{228}{\Sigma}
869 \DeclareInputMath{229}{\sigma}
870 \DeclareInputMath{230}{\mu}
871 \DeclareInputMath{231}{\gamma}
872 \DeclareInputMath{232}{\Phi}
873 \DeclareInputMath{233}{\theta}
874 \DeclareInputMath{234}{\Omega}
875 \DeclareInputMath{235}{\delta}
876 \DeclareInputMath{236}{\infty}
877 \DeclareInputMath{237}{\phi}
878 \DeclareInputMath{238}{\varepsilon}
879 \DeclareInputMath{239}{\cap}
880 \DeclareInputMath{240}{\equiv}
881 \DeclareInputMath{241}{\pm}
882 \DeclareInputMath{242}{\ge}
883 \DeclareInputMath{243}{\le}
884 %\DeclareInputMath{244}{}   % upper part of integral sign
885 %\DeclareInputMath{245}{}   % lower part of integral sign
886 \DeclareInputMath{246}{\div}
887 \DeclareInputMath{247}{\approx}
888 \DeclareInputText{248}{\textdegree}
889 \DeclareInputText{249}{\textperiodcentered}
890 \DeclareInputText{250}{\textbullet}
891 \DeclareInputMath{251}{\surd}
892 \DeclareInputMath{252}{\mathnsuperior}
893 \DeclareInputMath{253}{\mathtwosuperior}
894 \DeclareInputText{254}{\textblacksquare} % right name?
895 \DeclareInputText{255}{\nobreakspace}

896 ⟨/cp437 | cp437de | cp865⟩

13 The Macintosh encoding

This input encoding was contributed by Constantin Kahn (kahn@math.toronto.edu), with minor modifications by Alan Jeffrey.

897 ⟨∗applemac⟩

898 \DeclareInputText{128}{\"A}
899 \DeclareInputText{129}{\r A}
900 \DeclareInputText{130}{\c C}
901 \DeclareInputText{131}{\@tabacckludge'E}
902 \DeclareInputText{132}{\~N}
903 \DeclareInputText{133}{\"O}
904 \DeclareInputText{134}{\"U}
905 \DeclareInputText{135}{\@tabacckludge'a}
906 \DeclareInputText{136}{\@tabacckludge`a}
907 \DeclareInputText{137}{\^a}
908 \DeclareInputText{138}{\"a}
909 \DeclareInputText{139}{\~a}
910 \DeclareInputText{140}{\r a}
911 \DeclareInputText{141}{\c c}
912 \DeclareInputText{142}{\@tabacckludge'e}
913 \DeclareInputText{143}{\@tabacckludge`e}
914 \DeclareInputText{144}{\^e}
915 \DeclareInputText{145}{\"e}
916 \DeclareInputText{146}{\@tabacckludge'\i}
917 \DeclareInputText{147}{\@tabacckludge`\i}
918 \DeclareInputText{148}{\^\i}
919 \DeclareInputText{149}{\"\i}
920 \DeclareInputText{150}{\~n}
921 \DeclareInputText{151}{\@tabacckludge'o}
922 \DeclareInputText{152}{\@tabacckludge`o}
923 \DeclareInputText{153}{\^o}
924 \DeclareInputText{154}{\"o}
925 \DeclareInputText{155}{\~o}
926 \DeclareInputText{156}{\@tabacckludge'u}
927 \DeclareInputText{157}{\@tabacckludge`u}
928 \DeclareInputText{158}{\^u}
929 \DeclareInputText{159}{\"u}
930 \DeclareInputText{160}{\dag}
931 \DeclareInputText{161}{\textdegree}
932 \DeclareInputText{162}{\textcent}
933 \DeclareInputText{163}{\pounds}
934 \DeclareInputText{164}{\S}
935 \DeclareInputText{165}{\textbullet}
936 \DeclareInputText{166}{\P}
937 \DeclareInputText{167}{\ss}
938 \DeclareInputText{168}{\textregistered}
939 \DeclareInputText{169}{\copyright}
940 \DeclareInputText{170}{\texttrademark}
941 \DeclareInputText{171}{\@tabacckludge'{}}

942 \DeclareInputText{172}{\"{}}
943 \DeclareInputMath{173}{\neq}
944 \DeclareInputText{174}{\AE}
945 \DeclareInputText{175}{\O}
946 \DeclareInputMath{176}{\infty}
947 \DeclareInputMath{177}{\pm}
948 \DeclareInputMath{178}{\leq}
949 \DeclareInputMath{179}{\geq}
950 \DeclareInputText{180}{\textyen}
951 \DeclareInputMath{181}{\mu}
952 \DeclareInputMath{182}{\partial}
953 \DeclareInputMath{183}{\Sigma}
954 \DeclareInputMath{184}{\Pi}
955 \DeclareInputMath{185}{\pi}
956 \DeclareInputMath{186}{\int}
957 \DeclareInputMath{187}{\mathordfeminine}
958 \DeclareInputMath{188}{\mathordmasculine}
959 \DeclareInputMath{189}{\Omega}
960 \DeclareInputText{190}{\ae}
961 \DeclareInputText{191}{\o}
962 \DeclareInputText{192}{\textquestiondown}
963 \DeclareInputText{193}{\textexclamdown}
964 \DeclareInputMath{194}{\lnot}
965 \DeclareInputMath{195}{\surd}
966 \DeclareInputText{196}{\textflorin}
967 \DeclareInputMath{197}{\approx}
968 \DeclareInputMath{198}{\Delta}
969 \DeclareInputText{199}{\guillemotleft}
970 \DeclareInputText{200}{\guillemotright}
971 \DeclareInputText{201}{\dots}
972 \DeclareInputText{202}{\nobreakspace}
973 \DeclareInputText{203}{\@tabacckludge`A}
974 \DeclareInputText{204}{\~A}
975 \DeclareInputText{205}{\~O}
976 \DeclareInputText{206}{\OE}
977 \DeclareInputText{207}{\oe}
978 \DeclareInputText{208}{\textendash}
979 \DeclareInputText{209}{\textemdash}
980 \DeclareInputText{210}{\textquotedblleft}
981 \DeclareInputText{211}{\textquotedblright}
982 \DeclareInputText{212}{\textquoteleft}
983 \DeclareInputText{213}{\textquoteright}
984 \DeclareInputMath{214}{\div}
985 \DeclareInputMath{215}{\diamond}
986 \DeclareInputText{216}{\"y}
987 \DeclareInputText{217}{\"Y}
988 \DeclareInputMath{218}{/}
989 \DeclareInputText{219}{\textcurrency}
990 \DeclareInputText{220}{\guilsinglleft}

 991 \DeclareInputText{221}{\guilsinglright}
 992 \DeclareInputText{222}{fi}
 993 \DeclareInputText{223}{fl}
 994 \DeclareInputText{224}{\ddag}
 995 \DeclareInputText{225}{\textperiodcentered}
 996 \DeclareInputText{226}{\quotesinglbase}
 997 \DeclareInputText{227}{\quotedblbase}
 998 \DeclareInputText{228}{\textperthousand}
 999 \DeclareInputText{229}{\^A}
1000 \DeclareInputText{230}{\^E}
1001 \DeclareInputText{231}{\@tabacckludge'A}
1002 \DeclareInputText{232}{\"E}
1003 \DeclareInputText{233}{\@tabacckludge`E}
1004 \DeclareInputText{234}{\@tabacckludge'I}
1005 \DeclareInputText{235}{\^I}
1006 \DeclareInputText{236}{\"I}
1007 \DeclareInputText{237}{\@tabacckludge`I}
1008 \DeclareInputText{238}{\@tabacckludge'O}
1009 \DeclareInputText{239}{\^O}
1010 \DeclareInputText{240}{\textapplelogo}
1011 \DeclareInputText{241}{\@tabacckludge`O}
1012 \DeclareInputText{242}{\@tabacckludge'U}
1013 \DeclareInputText{243}{\^U}
1014 \DeclareInputText{244}{\@tabacckludge`U}
1015 \DeclareInputText{245}{\i}
1016 \DeclareInputText{246}{\^{}}
1017 \DeclareInputText{247}{\~{}}
1018 \DeclareInputText{248}{\@tabacckludge={}}
1019 \DeclareInputText{249}{\u{}}
1020 \DeclareInputText{250}{\.{}}
1021 \DeclareInputText{251}{\r{}}
1022 \DeclareInputText{252}{\c\ }
1023 \DeclareInputText{253}{\H{}}
1024 \DeclareInputText{254}{\k\ }
1025 \DeclareInputText{255}{\v{}}

1026 ⟨/applemac⟩

14 The Next encoding

This input encoding is based on work by Stefan Ried (stef@theo-phys.uni-essen.de

  and Holger Uhr (huhr@uni-paderborn.de).

1027 ⟨∗next⟩

1028 \DeclareInputText{128}{\nobreakspace}
1029 \DeclareInputText{129}{\@tabacckludge`A}
1030 \DeclareInputText{130}{\@tabacckludge'A}
1031 \DeclareInputText{131}{\^A}
1032 \DeclareInputText{132}{\~A}
1033 \DeclareInputText{133}{\"A}

1034 \DeclareInputText{134}{\r A}
1035 \DeclareInputText{135}{\c C}
1036 \DeclareInputText{136}{\@tabacckludge`E}
1037 \DeclareInputText{137}{\@tabacckludge'E}
1038 \DeclareInputText{138}{\^E}
1039 \DeclareInputText{139}{\"E}
1040 \DeclareInputText{140}{\@tabacckludge`I}
1041 \DeclareInputText{141}{\@tabacckludge'I}
1042 \DeclareInputText{142}{\^I}
1043 \DeclareInputText{143}{\"I}
1044 %\DeclareInputText{144}{?}
1045 \DeclareInputText{145}{\~N}
1046 \DeclareInputText{146}{\@tabacckludge`O}
1047 \DeclareInputText{147}{\@tabacckludge'O}
1048 \DeclareInputText{148}{\^O}
1049 \DeclareInputText{149}{\~O}
1050 \DeclareInputText{150}{\"O}
1051 \DeclareInputText{151}{\@tabacckludge`U}
1052 \DeclareInputText{152}{\@tabacckludge'U}
1053 \DeclareInputText{153}{\^U}
1054 \DeclareInputText{154}{\"U}
1055 \DeclareInputText{155}{\@tabacckludge'Y}
1056 %\DeclareInputText{156}{?}
1057 \DeclareInputMath{157}{\mu}
1058 \DeclareInputMath{158}{\times}
1059 \DeclarIenputMath{159}{\div}
1060 \DeclareInputText{160}{\copyright}
1061 \DeclareInputText{161}{\textexclamdown}
1062 %\DeclareInputText{162}{?}
1063 \DeclareInputText{163}{\pounds}
1064 \DeclareInputMath{164}{/}
1065 %\DeclareInputText{165}{?}
1066 \DeclareInputText{166}{\textflorin}
1067 \DeclareInputText{167}{\S}
1068 %\DeclareInputText{168}{?}
1069 \DeclareInputText{169}{\textquotesingle}
1070 \DeclareInputText{170}{\textquotedblleft}
1071 %\DeclareInputText{171}{?}
1072 %\DeclareInputText{172}{?}
1073 %\DeclareInputText{173}{?}
1074 \DeclareInputText{174}{fi}
1075 \DeclareInputText{175}{fl}
1076 %\DeclareInputText{176}{?}
1077 \DeclareInputText{177}{\textendash}
1078 \DeclareInputText{178}{\dag}
1079 \DeclareInputText{179}{\ddag}
1080 \DeclareInputText{180}{\textperiodcentered}
1081 \DeclareInputText{181}{\textbar}
1082 \DeclareInputText{182}{\P}

1083 \DeclareInputText{183}{\textbullet}
1084 \DeclareInputText{184}{\quotesinglbase}
1085 \DeclareInputText{185}{\quotedblbase}
1086 \DeclareInputText{186}{\textquotedblright}
1087 %\DeclareInputText{187}{?}
1088 \DeclareInputText{188}{\dots}
1089 %\DeclareInputText{189}{?}
1090 \DeclareInputMath{190}{\lnot}
1091 \DeclareInputText{191}{\textquestiondown}
1092 \DeclareInputMath{192}{\mathonesuperior}
1093 \DeclareInputText{193}{\@tabacckludge`{}}
1094 \DeclareInputText{194}{\@tabacckludge'{}}
1095 \DeclareInputText{195}{\^{}}
1096 \DeclareInputText{196}{\~{}}
1097 \DeclareInputText{197}{\@tabacckludge={}}
1098 \DeclareInputText{198}{\u{}}
1099 \DeclareInputText{199}{\.{}}
1100 \DeclareInputText{200}{\"{}}
1101 \DeclareInputMath{201}{\mathtwosuperior}
1102 \DeclareInputText{202}{\textdegree}
1103 \DeclareInputText{203}{\c\ }
1104 \DeclareInputMath{204}{\maththreesuperior}
1105 \DeclareInputText{205}{\H{}}
1106 %\DeclareInputText{206}{?}
1107 \DeclareInputText{207}{\v{}}
1108 \DeclareInputText{208}{\textemdash}
1109 \DeclareInputMath{209}{\pm}
1110 \DeclareInputText{210}{\textonequarter}
1111 \DeclareInputText{211}{\textonehalf}
1112 \DeclareInputText{212}{\textthreequarters}
1113 \DeclareInputText{213}{\@tabacckludge`a}
1114 \DeclareInputText{214}{\@tabacckludge'a}
1115 \DeclareInputText{215}{\^a}
1116 \DeclareInputText{216}{\~a}
1117 \DeclareInputText{217}{\"a}
1118 \DeclareInputText{218}{\r a}
1119 \DeclareInputText{219}{\c c}
1120 \DeclareInputText{220}{\@tabacckludge`e}
1121 \DeclareInputText{221}{\@tabacckludge'e}
1122 \DeclareInputText{222}{\^e}
1123 \DeclareInputText{223}{\"e}
1124 \DeclareInputText{224}{\@tabacckludge`\i}
1125 \DeclareInputText{225}{\AE}
1126 \DeclareInputText{226}{\@tabacckludge'\i}
1127 \DeclareInputMath{227}{\mathordfeminine}
1128 \DeclareInputText{228}{\^\i}
1129 \DeclareInputText{229}{\"\i}
1130 %\DeclareInputText{230}{?}
1131 \DeclareInputText{231}{\~n}

1132 \DeclareInputText{232}{\L}
1133 \DeclareInputText{233}{\O}
1134 \DeclareInputText{234}{\OE}
1135 \DeclareInputMath{235}{\mathordmasculine}
1136 \DeclareInputText{236}{\@tabacckludge`o}
1137 \DeclareInputText{237}{\@tabacckludge'o}
1138 \DeclareInputText{238}{\^o}
1139 \DeclareInputText{239}{\~o}
1140 \DeclareInputText{240}{\"o}
1141 \DeclareInputText{241}{\ae}
1142 \DeclareInputText{242}{\@tabacckludge`u}
1143 \DeclareInputText{243}{\@tabacckludge'u}
1144 \DeclareInputText{244}{\^u}
1145 \DeclareInputText{245}{\i}
1146 \DeclareInputText{246}{\"u}
1147 \DeclareInputText{247}{\@tabacckludge'y}
1148 \DeclareInputText{248}{\l}
1149 \DeclareInputText{249}{\o}
1150 \DeclareInputText{250}{\oe}
1151 \DeclareInputText{251}{\ss}
1152 %\DeclareInputText{252}{?}
1153 \DeclareInputText{253}{\"y}

1154 ⟨/next⟩

15 The MS Windows ANSI encoding cp1252, and the ISO Latin-1 encoding

The MS Windows ANSI, cp 1252, input encoding was contributed by Berthold

  K.P. Horn (bkph@ai.mit.edu).

It has two very different names so the same code produces two files. These encoding files both define the characters in the MS Windows 3.1 ANSI encoding (Western Europe), also known as code page 1252, which is based on ISO Latin-1 but has important additions in the 128–159 range.

Designed for: Danish, Dutch, English, Finnish, French, German, Icelandic, Italian, Norwegian, Portuguese, Spanish, and Swedish.

Note: Windows ANSI — like Macintosh standard roman encoding — has quotesingle at 39, and grave at 96 — which is here ignored.

The ISO Latin-1 encoding file defines only the restricted range of characters available in the ISO 8859-1 encoding.

1155 ⟨∗ansinew | cp1252⟩

1156 \DeclareInputText{130}{\quotesinglbase}
1157 \DeclareInputText{131}{\textflorin}
1158 \DeclareInputText{132}{\quotedblbase}
1159 \DeclareInputText{133}{\dots}
1160 \DeclareInputText{134}{\dag}
1161 \DeclareInputText{135}{\ddag}

1162 \DeclareInputText{136}{\^{}}
1163 \DeclareInputText{137}{\textperthousand}
1164 \DeclareInputText{138}{\v S}
1165 \DeclareInputText{139}{\guilsinglleft}
1166 \DeclareInputText{140}{\OE}
1167 \DeclareInputText{145}{\textquoteleft}
1168 \DeclareInputText{146}{\textquoteright}
1169 \DeclareInputText{147}{\textquotedblleft}
1170 \DeclareInputText{148}{\textquotedblright}
1171 \DeclareInputText{149}{\textbullet}
1172 \DeclareInputText{150}{\textendash}
1173 \DeclareInputText{151}{\textemdash}
1174 \DeclareInputText{152}{\~{}}
1175 \DeclareInputText{153}{\texttrademark}
1176 \DeclareInputText{154}{\v s}
1177 \DeclareInputText{155}{\guilsinglright}
1178 \DeclareInputText{156}{\oe}
1179 \DeclareInputText{159}{\"Y}

1180 ⟨/ansinew | cp1252⟩

1181 ⟨∗ansinew | cp1252 | latin1⟩

1182 \DeclareInputText{160}{\nobreakspace}
1183 \DeclareInputText{176}{\textdegree}
1184 \DeclareInputText{161}{\textexclamdown}
1185 \DeclareInputMath{177}{\pm}
1186 \DeclareInputText{162}{\textcent}
1187 \DeclareInputMath{178}{\mathtwosuperior}
1188 \DeclareInputText{163}{\pounds}
1189 \DeclareInputMath{179}{\maththreesuperior}
1190 \DeclareInputText{164}{\textcurrency}
1191 \DeclareInputText{180}{\@tabacckludge'{}}
1192 \DeclareInputText{165}{\textyen}
1193 \DeclareInputMath{181}{\mu}
1194 \DeclareInputText{166}{\textbrokenbar}
1195 \DeclareInputText{182}{\P}
1196 \DeclareInputText{167}{\S}
1197 \DeclareInputText{183}{\textperiodcentered}
1198 \DeclareInputText{168}{\"{}}
1199 \DeclareInputText{184}{\c\ }
1200 \DeclareInputText{169}{\copyright}
1201 \DeclareInputMath{185}{\mathonesuperior}
1202 \DeclareInputMath{170}{\mathordfeminine}
1203 \DeclareInputMath{186}{\mathordmasculine}
1204 \DeclareInputText{171}{\guillemotleft}
1205 \DeclareInputText{187}{\guillemotright}
1206 \DeclareInputMath{172}{\lnot}
1207 \DeclareInputText{188}{\textonequarter}
1208 \DeclareInputText{173}{\-}
1209 \DeclareInputText{189}{\textonehalf}
1210 \DeclareInputText{174}{\textregistered}

1211 \DeclareInputText{190}{\textthreequarters}
1212 \DeclareInputText{175}{\@tabacckludge={}}
1213 \DeclareInputText{191}{\textquestiondown}
1214 \DeclareInputText{192}{\@tabacckludge`A}
1215 \DeclareInputText{208}{\DH}
1216 \DeclareInputText{193}{\@tabacckludge'A}
1217 \DeclareInputText{209}{\~N}
1218 \DeclareInputText{194}{\^A}
1219 \DeclareInputText{210}{\@tabacckludge`O}
1220 \DeclareInputText{195}{\~A}
1221 \DeclareInputText{211}{\@tabacckludge'O}
1222 \DeclareInputText{196}{\"A}
1223 \DeclareInputText{212}{\^O}
1224 \DeclareInputText{197}{\r A}
1225 \DeclareInputText{213}{\~O}
1226 \DeclareInputText{198}{\AE}
1227 \DeclareInputText{214}{\"O}
1228 \DeclareInputText{199}{\c C}
1229 \DeclareInputMath{215}{\times}
1230 \DeclareInputText{200}{\@tabacckludge`E}
1231 \DeclareInputText{216}{\O}
1232 \DeclareInputText{201}{\@tabacckludge'E}
1233 \DeclareInputText{217}{\@tabacckludge`U}
1234 \DeclareInputText{202}{\^E}
1235 \DeclareInputText{218}{\@tabacckludge'U}
1236 \DeclareInputText{203}{\"E}
1237 \DeclareInputText{219}{\^U}
1238 \DeclareInputText{204}{\@tabacckludge`I}
1239 \DeclareInputText{220}{\"U}
1240 \DeclareInputText{205}{\@tabacckludge'I}
1241 \DeclareInputText{221}{\@tabacckludge'Y}
1242 \DeclareInputText{206}{\^I}
1243 \DeclareInputText{222}{\TH}
1244 \DeclareInputText{207}{\"I}
1245 \DeclareInputText{223}{\ss}
1246 \DeclareInputText{224}{\@tabacckludge`a}
1247 \DeclareInputText{240}{\dh}
1248 \DeclareInputText{225}{\@tabacckludge'a}
1249 \DeclareInputText{241}{\~n}
1250 \DeclareInputText{226}{\^a}
1251 \DeclareInputText{242}{\@tabacckludge`o}
1252 \DeclareInputText{227}{\~a}
1253 \DeclareInputText{243}{\@tabacckludge'o}
1254 \DeclareInputText{228}{\"a}
1255 \DeclareInputText{244}{\^o}
1256 \DeclareInputText{229}{\r a}
1257 \DeclareInputText{245}{\~o}
1258 \DeclareInputText{230}{\ae}
1259 \DeclareInputText{246}{\"o}

1260 \DeclareInputText{231}{\c c}
1261 \DeclareInputMath{247}{\div}
1262 \DeclareInputText{232}{\@tabacckludge`e}
1263 \DeclareInputText{248}{\o}
1264 \DeclareInputText{233}{\@tabacckludge'e}
1265 \DeclareInputText{249}{\@tabacckludge`u}
1266 \DeclareInputText{234}{\^e}
1267 \DeclareInputText{250}{\@tabacckludge'u}
1268 \DeclareInputText{235}{\"e}
1269 \DeclareInputText{251}{\^u}
1270 \DeclareInputText{236}{\@tabacckludge`\i}
1271 \DeclareInputText{252}{\"u}
1272 \DeclareInputText{237}{\@tabacckludge'\i}
1273 \DeclareInputText{253}{\@tabacckludge'y}
1274 \DeclareInputText{238}{\^\i}
1275 \DeclareInputText{254}{\th}
1276 \DeclareInputText{239}{\"\i}
1277 \DeclareInputText{255}{\"y}

1278 ⟨/ansinew | cp1252 | latin1⟩

16 The MS Windows encoding cp1250

This is an MS Windows encoding for Central and Eastern Europe, known as Code Page 1250; it was provided by Marcin Woliński wolinski@melkor.mimuw.edu.pl and edited by Chris Rowley (see v0.94 below) to make it consistent with other files.

1279 ⟨∗cp1250⟩

1280 %\DeclareInputText{128}{} % not used
1281 %\DeclareInputText{129}{} % n/u
1282 \DeclareInputText{130}{\quotesinglbase}
1283 %\DeclareInputText{131}{} % n/u
1284 \DeclareInputText{132}{\quotedblbase}
1285 \DeclareInputText{133}{\dots}
1286 \DeclareInputText{134}{\dag}
1287 \DeclareInputText{135}{\ddag}
1288 %\DeclareInputText{136}{} % n/u
1289 \DeclareInputText{137}{\textperthousand}
1290 \DeclareInputText{138}{\v S}
1291 \DeclareInputText{139}{\guilsinglleft}
1292 \DeclareInputText{140}{\@tabacckludge'S}
1293 \DeclareInputText{141}{\v T}
1294 \DeclareInputText{142}{\v Z}
1295 \DeclareInputText{143}{\@tabacckludge'Z}
1296 %\DeclareInputText{144}{} % n/u
1297 \DeclareInputText{145}{\textquoteleft}
1298 \DeclareInputText{146}{\textquoteright}
1299 \DeclareInputText{147}{\textquotedblleft}
1300 \DeclareInputText{148}{\textquotedblright}

1301 \DeclareInputText{149}{\textbullet}
1302 \DeclareInputText{150}{\textendash}
1303 \DeclareInputText{151}{\textemdash}
1304 %\DeclareInputText{152}{} % n/u
1305 \DeclareInputText{153}{\texttrademark}
1306 \DeclareInputText{154}{\v s}
1307 \DeclareInputText{155}{\guilsinglright}
1308 \DeclareInputText{156}{\@tabacckludge's}
1309 \DeclareInputText{157}{\v t} % t caron (t') ?
1310 \DeclareInputText{158}{\v z}
1311 \DeclareInputText{159}{\@tabacckludge'z}
1312 \DeclareInputText{160}{\nobreakspace}
1313 \DeclareInputText{161}{\v{}}
1314 \DeclareInputText{162}{\u{}}
1315 \DeclareInputText{163}{\L}
1316 \DeclareInputText{164}{\textcurrency}
1317 \DeclareInputText{165}{\k A}
1318 \DeclareInputText{166}{\textbrokenbar}
1319 \DeclareInputText{167}{\S}
1320 \DeclareInputText{168}{\"{}}
1321 \DeclareInputText{169}{\copyright}
1322 \DeclareInputText{170}{\c S}
1323 \DeclareInputText{171}{\guillemotleft}
1324 \DeclareInputMath{172}{\lnot}
1325 \DeclareInputText{173}{\-} % soft hyphen
1326 \DeclareInputText{174}{\textregistered}
1327 \DeclareInputText{175}{\.Z}
1328 \DeclareInputText{176}{\textdegree} % degree sign
1329 \DeclareInputMath{177}{\pm} % plus-minus
1330 \DeclareInputText{178}{\k\ }
1331 \DeclareInputText{179}{\l}
1332 \DeclareInputText{180}{\@tabacckludge'{}}
1333 \DeclareInputMath{181}{\mu} % micro sign
1334 \DeclareInputText{182}{\P}
1335 \DeclareInputText{183}{\textperiodcentered}
1336 \DeclareInputText{184}{\c\ }
1337 \DeclareInputText{185}{\k a}
1338 \DeclareInputText{186}{\c s}
1339 \DeclareInputText{187}{\guillemotright}
1340 \DeclareInputText{188}{\v L} % L caron (L') ?
1341 \DeclareInputText{189}{\H{}}
1342 \DeclareInputText{190}{\v l} % l caron (l') ?
1343 \DeclareInputText{191}{\. z}
1344 \DeclareInputText{192}{\@tabacckludge'R}
1345 \DeclareInputText{193}{\@tabacckludge'A}
1346 \DeclareInputText{194}{\^A}
1347 \DeclareInputText{195}{\u A}
1348 \DeclareInputText{196}{\"A}
1349 \DeclareInputText{197}{\@tabacckludge'L}

1350 \DeclareInputText{198}{\@tabacckludge'C}
1351 \DeclareInputText{199}{\c C}
1352 \DeclareInputText{200}{\v C}
1353 \DeclareInputText{201}{\@tabacckludge'E}
1354 \DeclareInputText{202}{\k E}
1355 \DeclareInputText{203}{\"E}
1356 \DeclareInputText{204}{\v E}
1357 \DeclareInputText{205}{\@tabacckludge'I}
1358 \DeclareInputText{206}{\^I}
1359 \DeclareInputText{207}{\v D}
1360 \DeclareInputText{208}{\DJ} % D stroke
1361 \DeclareInputText{209}{\@tabacckludge'N}
1362 \DeclareInputText{210}{\v N}
1363 \DeclareInputText{211}{\@tabacckludge'O}
1364 \DeclareInputText{212}{\^O}
1365 \DeclareInputText{213}{\H O}
1366 \DeclareInputText{214}{\"O}
1367 \DeclareInputMath{215}{\times}
1368 \DeclareInputText{216}{\v R}
1369 \DeclareInputText{217}{\r U}
1370 \DeclareInputText{218}{\@tabacckludge'U}
1371 \DeclareInputText{219}{\H U}
1372 \DeclareInputText{220}{\" U}
1373 \DeclareInputText{221}{\@tabacckludge'Y}
1374 \DeclareInputText{222}{\c T}
1375 \DeclareInputText{223}{\ss}
1376 \DeclareInputText{224}{\@tabacckludge'r}
1377 \DeclareInputText{225}{\@tabacckludge'a}
1378 \DeclareInputText{226}{\^a}
1379 \DeclareInputText{227}{\u a}
1380 \DeclareInputText{228}{\" a}
1381 \DeclareInputText{229}{\@tabacckludge'l}
1382 \DeclareInputText{230}{\@tabacckludge'c}
1383 \DeclareInputText{231}{\c c}
1384 \DeclareInputText{232}{\v c}
1385 \DeclareInputText{233}{\@tabacckludge'e}
1386 \DeclareInputText{234}{\k e}
1387 \DeclareInputText{235}{\"e}
1388 \DeclareInputText{236}{\v e}
1389 \DeclareInputText{237}{\@tabacckludge'\i}
1390 \DeclareInputText{238}{\^\i}
1391 \DeclareInputText{239}{\v d} % d caron (d') ?
1392 \DeclareInputText{240}{\dj} % d stroke
1393 \DeclareInputText{241}{\@tabacckludge'n}
1394 \DeclareInputText{242}{\v n}
1395 \DeclareInputText{243}{\@tabacckludge'o}
1396 \DeclareInputText{244}{\^o}
1397 \DeclareInputText{245}{\H o}
1398 \DeclareInputText{246}{\"o}

1399 \DeclareInputMath{247}{\div}
1400 \DeclareInputText{248}{\v r}
1401 \DeclareInputText{249}{\r u}
1402 \DeclareInputText{250}{\@tabacckludge'u}
1403 \DeclareInputText{251}{\H u}
1404 \DeclareInputText{252}{\"u}
1405 \DeclareInputText{253}{\@tabacckludge'y}
1406 \DeclareInputText{254}{\c t}
1407 \DeclareInputText{255}{\.{}}

1408 ⟨/cp1250⟩

Finally, we reset the category code of the at sign at the end of all .def files.

1409 ⟨−package⟩\makeatother

Change History

v0.01

General: Created file. . . . . . . . . . . 1 v0.02

General: Added \ensuremath to some math commands. . . . . . . 1

Added \inputencoding. . . . . . . 1

Replaced \Dh by \DH and \Th by \TH. . . . . . . . . . . . . . . . . . . . . 1

v0.03

General: Added \DeclareInputComposite
and the \ProvidesCommands to

the encoding files. . . . . . . . . . . 1 Removed the definition of the accent slots in Latin-1. . . . . . . . . 1

v0.04

General:  Improved  coding

of  \DeclareInputText

and changed name from

\DeclareInputCharacter. . . . . 3

Made many Latin-1 characters math-only. . . . . . . . . . . . . . . . 1

Removed \DeclareInputComposite.

. . . . . . . . . . . . . . . . . . . . . . . . 1

Replaced \DeclareInputCharacter
by \DeclareInputText and
\DeclareInputMath. . . . . . . . . 1

v0.05

General: Updated for the new ver-

sion of ltoutenc. . . . . . . . . . . . 1 v0.06

General: Added \textregistered. 1

Added slot hex A0 to Latin-1. . 1 v0.07

General: Fixed typo A1 rather than A0. . . . . . . . . . . . . . . . . . . . . 1 Fixed where docstrip option had moved a line. . . . . . . . . . . . . . 1

v0.09

General: Added ‘beta test’ message. . . . . . . . . . . . . . . . . . . . 1 Made letters active and undefined by default, rather than illegal. . . . . . . . . . . . . . . . . . . . 1

v0.09c

General: Made mac encoding ‘active german quote safe’, and added the correct docstrip magic. . . . . . . . . . . . . . . . . . 24

v0.91

General:  Exchanged  codes

for  \textendash  and

\textemdash. . . . . . . . . . . . . 29

Put \makeatletter. . . \makeatother

around all .def files. . . . . . . . . . 1 Replaced \dh/\DH by \dj/\DJ. . . . . . . . . . . . . . . . . . . . . . 7, 18

v0.92

General: Added decmulti.def provided by M.Y. Chartoire. pr/2599 . . . . . . . . . . . . . . . . . 1

v0.93

General: \textperthousand not

\textpermill. pr/2673 . . . . . . 1 v0.94

General: Added cp1252 and merged latin1 . . . . . . . . . . . . . . . . . . 29 Added to and tidied documentation . . . . . . . . . . . . . . . . . . . . 1

Changed 158 to \DeclareInputText

. . . . . . . . . . . . . . . . . . . . . . . 21 Changed 176 to \textdegree . . 7 Changed \ldots to \dots, this should be undetectable since \ldots is not robust. . . . . . . . 29

Changed \textdagger to \dag

for consistency. . . . . . . . . . . . 32

Changed \textellipsis to

\dots for consistency. … 26, 32

Changed \textparagraph to \P

for consistency. . . . . . . . . . . . 32

Changed \textquotedblbase to
\quotedblbase: this may be

temporary. . . . . . . . . . . . . . . 26

Changed \textquotesinglbase
to \quotesinglbase: this may

be temporary. . . . . . . . . . . . . 26

Changed \textsection to \S for

consistency. . . . . . . . . . . . . . 32 Made degree a text glyph . . . . . 1 Merged latin1 with ansinew/cp1252 . . . . . . . . . . . . . . . . . . . . . . . . 7 NOTE: for consistency, when available the robust text-or-math internal form is always used – THIS MAY CHANGE . 1 Ogonek: changed \k{} to \k\ . . 1 Removed entries that were solely in next.def . . . . . . . . . . . . . . . 5

\IeC: Changed non-typeset case

from \protect to \noexpand:

temporary fix. . . . . . . . . . . . . . 4

\inputencoding: Changed to work

only in outer vmode, see latex/2608 . . . . . . . . . . . . . . . . . 4 Warning messaage reworded and line number added, also for latex/2608 . . . . . . . . . . . . . . . . . 4

v0.95 General: Updated documentation . 1

\inputencoding: Changed to work

in any vmode, for David . . . . . 4 v0.9a

                                   General: \textonequarter and

friends should be declared with

                                       \DeclareInputText . . . . . . . 29

Default settings moved to own section . . . . . . . . . . . . . . . . . . 5

v0.9b

General: Added Mac encoding, applemac.def . . . . . . . . . . . . . . . 1 Corrected ordmasc and ordfem which had been switched by mistake. . . . . . . . . . . . . . . . . . 5

\IeC: Added hackery with \IeC in order to avoid space being gobbled. . . . . . . . . . . . . . . . . . . . 3

v0.9d

General: Added cp437 . . . . . . . . . 1 Made changes to cp850 suggested by Christian Bartels . . . . . . . 16

v0.9e

General: Added a 0 to all character codes . . . . . . . . . . . . . . . . . . 21 Redeclared ”AD to be soft hyphen. . . . . . . . . . . . . . . . . 7, 29 Swapped ordfeminine and ordmasculine. . . . . . . . . . . . . . . 29

v0.9f

General: Added Next encoding,

next.def . . . . . . . . . . . . . . . . . 1 v0.9g

General: Replaced \' \` \= by \a' \a` \a= in order to get correct accents in a tabbing environment. . . . . . . . . . . . . . . . . . . . 3

v0.9h

General: Added Windows 3.1 ANSI

encoding, ansinew.def . . . . . . . 1 v0.9i

General: Changed internal name \a

                                       to \@tabacckludge to protect

against redefinition by malicicous users. . . . . . . . . . . . . . . . 1 Wrapped long lines . . . . . . . . . 1

v0.9j

General: Remove \endinput so doc-

strip reaches ansinew encoding 1 v0.9k

General: Added default commands for Next input encoding. . . . . . 5 Added error messages for unavailable characters. . . . . . . . . 6

Moved \textregistered and
\texttrademark to the kernel. . 5

Replaced uses of \textsterling with \pounds. . . . . . . . . . . . . . 1 Tidied up this encoding. . . . . . 26

v0.9l

General: Removed extraneous braces from the ‘superior’ glyphs. . . . . . . . . . . . . . . . . . . 7

v0.9m

General: Added \@tabacckludge

hacks. . . . . . . . . . . . . . . . . . . 1 Added \ensuremath to definitions of the fraction glyphs. … 5 Added German version . . . . . . . 1 Made bullet and periodcentered text glyphs rather than math glyphs. . . . . . . . . . . . . . . . . . . 1 Made fraction a math character. 26 Made uumlaut and pounds text characters . . . . . . . . . . . . . . . 21

Replaced \textasciitilde by
\nobreakspace. . . . . . . . . . . . 1

\IeC: Modified \IeC in order

to work in \hyphenation la-

tex/2004. . . . . . . . . . . . . . . . . 4 v0.9n

General: Made fraction a text char-

acter. . . . . . . . . . . . . . . . . . . 26 v0.9o

General: 0B2 corrected in latin2 latex/2079 . . . . . . . . . . . . . . . . . 1 0F0 corrected in cp850 latex/2080 . . . . . . . . . . . . . . . . . 1 Cedilla (\c\) rather than (\c{}) latex/2077, finished on 1996/10/28 . . . . . . . . . . . . . . . 1

v0.9p

General: ansinew 09f is Ÿ not ÿ, la-

tex/2119 . . . . . . . . . . . . . . . . 29 v0.9q

\@inpenc@loop: Macro added …. 5 \inputencoding: Allow characters

below 32 for latex/2071 . . . . . . 4 Check the def file was reasonable for latex/2136 . . . . . . . . . . . . . 4 v0.9r

General: Added cp852.def . . . . . . 1 v0.9s

General: Swapped mathordfemi-

nine and masculine. /2203 … 16 v0.9t

                                   General: Added \textblacksquare 5
                                      Added  \textflorin  and

\textpeseta . . . . . . . . . . . . . 21 Added cp865.def . . . . . . . . . . . 1

                                      Added extra \nobreakspace:

OK? . . . . . . . . . . . . . . . . . . . 18 Changed \aa and \AA to \r a and \r A . . . . . . . . . . . . . . . . . . . 1

                                      Changed \emptyset to \phi . . 21
                                      Changed \maththreesuperior
                                       to \mathnsuperior . . . . . . . . 21

Changed \Pi to \pi, perhaps . 21 Changed \Theta to \Phi …. 21

                                      Removed \textbrokenbar . . . 21

Removed \textendash . . . . . . 21 v0.9u

General: Added cp865 and corrected cp850, cp852 and cp437 guards . . . . . . . . . . . . . . . . . . 5 Added more to cp865.def . . . . . 1

                                      Corrected code for \textblacksquare

. . . . . . . . . . . . . . . . . . . . . . . . 5 v0.9v

\@inpenc@loop: Exit from the top

of loop fixed for latex/2257 … 5 v0.9w

General: Correct documentation of

                                       \@tabacckludge . . . . . . . . . . . 3

v0.9x

General: Use decimal rather than hex to avoid active character problems. latex/2451. . . . . . . . 3

v0.9y

General: Added latin5.def (provided by H. Turgut Uyar: uyar@cs.itu.edu.tr) . . . . . . . . . 1

v0.9z

General: Added latin3.def (provided by Jörg Knappen and modified by Chris Rowley) … 1 Corrected typo in slot 213 … 18

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

Symbols \" … 186, 210, 215, 224, 227, 242, 247, 256, 259, 285, 309, 314, 323, 326, 331, 341, 346, 355, 358, 363, 383, 407, 412, 421, 424, 429, 439, 444, 453, 456, 461, 462, 494, 499, 508, 511, 513, 515, 524, 529, 538, 541, 543, 545, 550, 556, 557, 565– 567, 569, 570, 576, 606, 613, 638, 655, 661, 662, 671, 672, 674, 681, 713, 740, 756, 759, 764, 766, 769, 775, 779–781, 898, 903, 904, 908, 915, 919, 924, 929, 942, 986, 987, 1002, 1006,  1033, 1039,  1043, 1050,  1054, 1100,  1117, 1123,  1129, 1140,  1146, 1153,  1179, 1198,  1222, 1227,  1236, 1239,  1244, 1254,  1259, 1268,  1271, 1276,  1277, 1320,  1348,

1355, 1366, 1372, 1380, 1387, 1398, 1404

\- . . . . . . . . . . 196,

295, 393, 1208, 1325 \. . . . . . . . 200, 201,

265, 287, 299, 300, 311, 312, 343, 344, 364, 426, 705, 706, 742, 1020, 1099, 1327, 1343, 1407

\@depth . . . . . . . . . . 105 \@empty . . . . . . . . . . 37

\@firstofone . . . . . . 46

\@height . . . . . . . . . 105

\@inpenc@loop 54–57, \underline{66}
\@inpenc@test 25, 52, 60

\@ne . . . . . . . . . . . . . 77

\@tabacckludge 179,

182, 183, 194, 195, 202, 204, 205, 209, 212, 214, 220, 223, 228, 229, 234, 236, 237, 241, 244, 246, 252, 255, 260, 261, 277, 301, 303, 306, 308, 317, 319, 320, 322, 325, 327, 333, 335, 338, 340, 349, 351, 352, 354, 357, 359, 376, 397, 399, 401, 404, 406, 415, 417, 418, 420, 423, 425, 431, 433, 436, 438, 447, 449, 450, 452, 455, 457, 487, 488,

491, 493, 502, 504, 505, 507, 510, 512, 517, 518, 521, 523, 532, 534, 535, 537, 540, 542, 549, 552, 558, 559, 563, 568, 574, 580–583, 586, 590, 607, 610, 617, 619, 625, 637, 641, 643, 645, 647, 649, 654, 656– 658, 665, 668, 670, 679, 683, 685–688, 696, 701, 714, 716, 721, 727, 729, 737, 739, 741, 745, 747, 751, 757, 760, 765, 768, 771, 776, 778, 796–799, 901, 905, 906, 912, 913, 916, 917, 921, 922, 926, 927, 941, 973, 1001, 1003, 1004,  1007, 1008,  1011, 1012,  1014, 1018,  1029, 1030,  1036, 1037,  1040, 1041,  1046, 1047,  1051, 1052,  1055, 1093,  1094, 1097,  1113, 1114,  1120, 1121,  1124, 1126,  1136,

1137, 1142, 1143, 1147, 1191, 1212, 1214, 1216, 1219, 1221, 1230, 1232, 1233, 1235, 1238, 1240, 1241, 1246, 1248, 1251, 1253, 1262, 1264, 1265, 1267, 1270, 1272, 1273, 1292, 1295, 1308, 1311, 1332, 1344, 1345, 1349, 1350, 1353, 1357, 1361, 1363, 1370, 1373, 1376, 1377, 1381, 1382, 1385, 1389, 1393, 1395, 1402, 1405

\@tempa . . . . . 34, 36, 37 \@tempb . . . . . . . . 35, 36

\@tempcnta . . . . . . .

. 67, 69, 71, 76, 77

\@typeset@protect . 45
\@undefined . . . . . . . 74

\@width . . . . . . . . . . 105 \^. . . . . . . . . 54–57,

206, 211, 230, 238, 243, 262, 280–282, 293, 294, 305, 310, 313, 318, 321, 324, 329, 330, 337, 342, 345, 350, 353, 356, 361, 362, 403, 408, 419, 422, 427, 435, 440, 451, 454, 459, 490, 495, 506, 509, 514, 520, 525, 536, 539, 544, 554, 555,

561, 564, 572, 588, 605, 612, 623, 639, 659, 660, 677, 702, 717, 725, 758, 763, 767, 774, 777, 907, 914, 918, 923, 928, 999, 1000, 1005, 1009,  1013, 1016,  1031, 1038,  1042, 1048,  1053, 1095,  1115, 1122,  1128, 1138,  1144, 1162,  1218, 1223,  1234, 1237,  1242, 1250,  1255, 1266,  1269, 1274,  1346, 1358,  1364, 1378, 1390, 1396 \~ . . . . . . . . . 27, 71, 304, 336, 402, 405, 410, 434, 437, 442, 489, 492, 497, 519, 522, 527, 584, 585, 609, 611, 627, 629, 800, 801, 902, 909, 920, 925, 974, 975, 1017, 1032, 1045,  1049, 1096,  1116, 1131,  1139, 1174,  1217, 1220,  1225, 1249, 1252, 1257 \␣ . . . . . . . 175, 187, 286, 384, 634, 726, 736, 1022, 1024,  1103, 1199, 1330, 1336

A \active . . . . . . . . . . 69 \advance . . . . . . . . . 77 \AE . . 411, 498, 553,

773, 944, 1125, 1226 \ae . . 443, 528, 551,

772, 960, 1141, 1258 \alpha . . . . . . . . . . . 861 \approx . . . . . . 887, 967

B \beta . . . . . . . . . . . . 863 \bgroup . . . . . . . . 26, 70

C \c 187, 190, 191, 216,

231, 248, 263, 286, 289, 290, 315, 347, 384, 413, 428, 445, 460, 500, 530, 548, 562, 634, 653, 667, 698, 704, 719, 736, 749, 755, 762, 900, 911, 1022, 1035, 1103, 1119, 1199, 1228, 1260, 1322, 1336, 1338, 1351, 1374, 1383, 1406

\cap . . . . . . . . . . . . . 879 \catcode . . . . . . . . . 69 \circ . . . . . . . . . . 83, 86

                                             \copyright . . . 385,

478, 592, 939, 1060, 1200, 1321

                                             \CurrentOption . . . . 79

D \dag 930, 1078, 1160, 1286 \ddag 994, 1079, 1161, 1287

                                             \DeclareInputMath .

. . \underline{23}, 217, 249, 273, 275, 279, 316, 348, 370, 372, 374, 378, 386–388, 391, 414, 446, 467, 469, 471, 473, 479–481, 577, 587, 589, 594, 622, 631, 632, 642, 644, 646, 682, 734, 802, 803, 806, 816, 822, 823, 825, 828, 829, 845, 854, 855, 861, 863, 866–887, 891–893, 943, 946–949, 951– 959, 964, 965, 967, 968, 984, 985, 988, 1057, 1058, 1064, 1090, 1092, 1101, 1104, 1109, 1127, 1135, 1185, 1187, 1189, 1193, 1201– 1203, 1206, 1229, 1261, 1324, 1329, 1333, 1367, 1399

\DeclareInputText .

. . . . . \underline{23}, 170– 216, 218–248, 250–265, 268– 272, 274, 276– 278, 280–315, 317–347, 349– 364, 367–369, 371, 373, 375– 377, 379–385, 389, 390, 392– 413, 415–445, 447–462, 465, 466, 468, 470, 472, 474–478, 482–545, 548– 576, 578–586, 588, 590–593, 595–601, 603– 613, 616–621, 623, 625–630, 633–641, 643, 645, 647–650, 653–681, 683– 733, 735–752, 755–781, 784– 786, 789–791, 794–801, 804, 805, 807–815, 817–821, 824, 826, 827, 830– 844, 846–853, 856–860, 864, 888–890, 894, 895, 898–942, 944, 945, 950, 960–963, 966, 969–983, 986, 987, 989–1025, 1028–1056, 1060–1063, 1065–1089, 1091, 1093– 1100, 1102, 1103, 1105– 1108, 1110– 1126, 1128– 1134, 1136– 1153, 1156– 1179, 1182– 1184, 1186, 1188, 1190– 1192, 1194– 1200, 1204, 1205, 1207– 1228, 1230– 1260, 1262– 1277, 1280– 1323, 1325– 1328, 1330– 1332, 1334– 1366, 1368– 1398, 1400–1407

                   \DeclareOption . . . . 79
                   \DeclareRobustCommand

. . . . . . . . . . . . 43

                   \DeclarIenputMath 1059

\def . . . . . . . . 24, 30,

33–35, 44, 51, 66 \Delta . . . . . . . . . . . 968 \delta . . . . . . . . . . . 875 \DH . . . . . . . . 604, 1215 \dh . . . . . . . . 603, 1247 \diamond . . . . . . . . . 985 \div 249, 348, 446, 632,

734, 886, 984, 1059, 1261, 1399

\DJ . . . . . 203, 711, 1360 \dj . . . . . 235, 710, 1392 \dots 971, 1088, 1159, 1285

E \edef . . . . . . . . . . . . 36 \egroup . . . . . . . . 29, 73 \else . . . . . . . 39, 47, 60

                                             \ensuremath . . . . 83,

86, 89, 92, 95, 828, 829, 854, 855

\equiv . . . . . . . . . . . 880

                                             \expandafter . . . . 36, 46

F \fi . . . . . 41, 49, 59, 64 \frac . . . . . . . 89, 92, 95

G \Gamma . . . . . . . . . . . 866 \gamma . . . . . . . . . . . 871 \gdef . . . . . . . . . . . . 52 \ge . . . . . . . . . . . . . 882 \geq . . . . . . . . . . . . . 949 \global . . . . . . . . . . 52

                                             \guillemotleft 389,

482, 599, 699, 810, 969, 1204, 1323

                                             \guillemotright 390,

483, 601, 700, 811, 970, 1205, 1339

                                             \guilsinglleft . . . .

. . 990, 1165, 1291

                                             \guilsinglright . . .

. . 991, 1177, 1307

H \H … 197, 213, 225,

245, 257, 673, 675, 724, 743, 744, 1023, 1105, 1341, 1365, 1371, 1397, 1403

I \i 260, 262, 288, 357,

359, 361, 363, 455, 457–459, 461, 540, 542, 544, 545, 570, 572, 574, 581, 608, 677, 686, 766–768, 797, 916–919, 1015, 1124, 1126, 1128, 1129, 1145, 1270, 1272, 1274, 1276, 1389, 1390

\IeC . . . . . . . . . . . . . \underline{23} \ifnum . . . . . . . . . . . 76 \ifvmode . . . . . . . . . 53 \ifx . . . . . . . . 37, 45, 60 \infty . . . . . . . 876, 946 \input . . . . . . . . . . . 58

\inputencoding . . \underline{51}, 79

\int . . . . . . . . . . . . . 956

J

\j . . . . . . . . . . . . . . 294

K

\k . . . . . . . 172, 173, 175, 222, 254, 689, 690, 693, 694, 726, 1024, 1317, 1330, 1337, 1354, 1386

L

\L . 176, 680, 1132, 1315 \l . 177, 669, 1148, 1331 \lceil . . . . . . . . . . . 855 \le . . . . . . . . . . . . . 883 \leq . . . . . . . . . . . . . 948 \let . . . . . . . . . . . 52, 74 \lfloor . . . . . . . . . . 829 \llap . . . . . . . . . . . . 281 \lnot . . . . . . . . 391,

594, 806, 964, 1090, 1206, 1324

\loop . . . . . . . . . . . . 68

M

\makeatletter . . . . . 22
\makeatother . . . . . 1409
\mathnsuperior 158, 892

                          \mathonesuperior . .

…. 146, 386, 479, 642, 1092, 1201

                          \mathordfeminine . .

…. 163, 167, 387, 480, 587, 802, 957, 1127, 1202

                          \mathordmasculine .

…. 162, 166, 388, 481, 589, 803, 958, 1135, 1203

                          \maththreesuperior .

. 149, 275, 374, 471, 644, 1104, 1189

                          \mathtwosuperior . .

…. 152, 155, 273, 372, 469, 646, 893, 1101, 1187

\meaning . . . . . . . . . 36 \message . . . . . . . . . . 2

                          \MessageBreak . . . 62, 63

\mu . . 279, 378, 473, 631, 870, 951, 1057, 1193, 1333

N

                          \NeedsTeXFormat . . . . 1

\neq . . . . . . . . . . . . . 943

                          \nobreakspace . . . . .

…. 170, 268, 367, 650, 695, 752, 895, 972, 1028, 1182, 1312

\noexpand . . . . . . . . 48 \notdef 278, 297, 298,

302, 307, 334, 339

O \O … 416, 503, 575,

791, 945, 1133, 1231 \o … 448, 533, 571,

789, 961, 1149, 1263 \OE 501, 976, 1134, 1166 \oe 531, 977, 1150, 1178 \Omega . . . . . . . 874, 959

P \P 380, 474, 628, 936,

1082, 1195, 1334

                          \PackageWarning . . . 61

\partial . . . . . . . . . 952 \Phi . . . . . . . . . . . . . 872 \phi . . . . . . . . . . . . . 877 \Pi . . . . . . . . . . . . . 954 \pi . . . . . . . . . 867, 955 \pm . . . . . . 370, 467,

622, 881, 947, 1109, 1185, 1329

\pounds … 274, 373, 470, 573, 785, 790, 933, 1063, 1188

                                                   \ProcessOptions . . . 80

\protect . . . . . . . . . 45

                                                   \providecommand . . .

…. 146, 149, 152, 155, 158, 162, 163, 166, 167

                                                   \ProvidesFile . . . . 4–19
                                                   \ProvidesPackage . . . 3
                                                   \ProvideTextCommandDefault

. . . . . . 83, 86, 89, 92, 95, 98, 101, 104, 108, 110, 114, 116, 120, 124, 128, 132, 134, 138, 142

Q

                                                   \quotedblbase . 997,

1085, 1158, 1284

                                                   \quotesinglbase 996,

1084, 1156, 1282

R \r 221, 253, 409, 441,

496, 526, 560, 578, 663, 720, 761, 770, 899, 910, 1021, 1034, 1118, 1224, 1256, 1369, 1401

\rceil . . . . . . . . . . . 828 \relax 52, 60, 67, 76, 105 \repeat . . . . . . . . . . 78 \rfloor . . . . . . . . . . 854

S \S 184, 283, 381, 475,

630, 732, 934, 1067, 1196, 1319

\Sigma . . . . . . . 868, 953 \sigma . . . . . . . . . . . 869 \ss . . . . . . 233, 332,

430, 516, 621, 723, 864, 937, 1151, 1245, 1375

\surd . . . . . . . . 891, 965

T

\textapplelogo . . . 1010

\textbar . . . . . . . . 1081

\textblacksquare . .

. 104, 648, 750, 894

\textbrokenbar . . . .

…. 128, 129, 379, 616, 1194, 1318

\textbullet 890, 935,

1083, 1171, 1301

\textcent . . . . . . . .

. 108, 109, 114, 115, 371, 468, 598, 784, 932, 1186

\textcurrency . 120,

121, 124, 125, 178, 276, 375, 477, 618, 709, 812, 989, 1190, 1316

\textdarkgraybox . . 815
\textdegree . . . . . . .

83, 86, 171, 269, 368, 465, 636, 738, 888, 931, 1102, 1183, 1328

\textemdash . . . 979,

1108, 1173, 1303

\textendash . . . . . . .

. 620, 722, 978, 1077, 1172, 1302

\textexclamdown . . .

. 369, 466, 597, 809, 963, 1061, 1184

\textflorin . 98, 579,

795, 966, 1066, 1157

\textgraybox . . . . . . 814

\textit . . . . . . . . . . 98

\textlightgraybox . 813

\textmalteseH . . . . .

…. 132, 133, 270

\textmalteseh . . . . .

…. 134, 135, 271

\textonehalf . . . 89,

296, 394, 485, 595, 807, 1111, 1209

\textonequarter . . .

. . 92, 392, 484, 596, 808, 1110, 1207

\textperiodcentered

. 284, 382, 476, 640, 889, 995, 1080, 1197, 1335

\textperthousand 142,

143, 998, 1163, 1289

\textpeseta . . . 101, 794
\textquestiondown .

. 398, 486, 591, 804, 962, 1091, 1213

\textquotedblleft .

. . . . . . . . 980, 1070, 1169, 1299

\textquotedblright .

. . . . . . . . 981, 1086, 1170, 1300

\textquoteleft . . . .

. . 982, 1167, 1297

\textquoteright . . .

. . 983, 1168, 1298

\textquotesingle . .

… 138, 139, 1069

\textregistered 395,

593, 938, 1210, 1326

\TextSymbolUnavailable

. . . . . . . . 109, 111, 115, 117, 121, 125, 129, 133, 135, 139, 143

\textthreequarters .

. . . . . . . . . 95, 396, 626, 1112, 1211

\texttrademark . . . .

. . 940, 1175, 1305 \textyen 110, 111, 116,

117, 377, 472, 600, 786, 950, 1192

\TH . . . . . . . . 635, 1243 \th . . . . . . . . 633, 1275 \theta . . . . . . . . . . . 873 \times …. 217, 316,

414, 577, 682, 1058, 1229, 1367

U

\u … 174, 208, 240, 272, 291, 292, 328, 360, 400, 432, 707, 708, 730, 1019, 1098, 1314, 1347, 1379

\uccode . . . . . . . . 27, 71

\uppercase . . . . . 28, 72

V

\v … 180, 181, 185, 188, 189, 192, 193, 198, 199, 207, 218, 219, 226, 232, 239, 250, 251, 258, 264, 664, 666, 676, 678, 684, 691, 692, 697, 703, 712, 715, 718, 728, 731, 733, 735, 746, 748, 1025, 1107, 1164, 1176, 1290, 1293, 1294, 1306, 1309, 1310, 1313, 1340, 1342, 1352, 1356, 1359, 1362, 1368, 1384, 1388, 1391, 1394, 1400

\varepsilon . . . . . . . 878

\vrule . . . . . . . . . . . 105